### v1

# https://sinhasaptarshi.github.io/escounts

Abstract
Video repetition counting infers the number of repetitions of recurring actions or motion within a video. We propose an exemplar-based approach that discovers visual correspondence of video exemplars across repetitions within target videos. Our proposed Every Shot Counts (ESCounts) model is an attention-based encoder-decoder that encodes videos of varying lengths alongside exemplars from the same and different videos. In training, ESCounts regresses locations of high correspondence to the exemplars within the video. In tandem, our method learns a latent that encodes representations of general repetitive motions, which we use for exemplar-free, zero-shot inference. Extensive experiments over commonly used datasets (RepCount, Countix, and UCFRep) showcase ESCounts obtaining state-of-the-art performance across all three datasets. On RepCount, ESCounts increases the off-by-one from $0.39$ to $0.56$ and decreases the mean absolute error from $0.38$ to $0.21$. Detailed ablations further demonstrate the effectiveness of our method.
Keywords:
Video Repetition Counting Video Exemplar Cross-Attention Transformer Video Understanding
1 Introduction
Recent years have come with tremendous progress in video understanding. Large Visual Language Models (VLMs) have been adopted for many vision tasks including video summarisation [44, 32, 60], localisation [45, 57], and question answering (VQA) [1, 21, 39, 64]. Despite their great success, recent analysis [24] shows that VLMs can still fail to correctly count objects or actions. Robust counting can be challenging due to appearance diversity, limited training data, and most critically semantic ambiguity – identifying ‘what’ exactly to count.
Evidence in developmental psychology and cognitive neuroscience [50, 53, 54] shows that infants fail to differentiate the number of hidden objects when they are not shown and counted to them first, suggesting an upper limit of individual objects that can be represented in working memory. However, infants that were first exposed to an instance of the object could better approximate the carnality. This shows that counting is a visual exercise of matching to given exemplars, and is developed in humans prior to understanding their semantics.
Object-counting in images has recently exploited exemplars to improve performance [38, 10]. In training, models attend to one or more exemplars of ‘what’ object(s) to count alongside learnt embeddings for exemplar-free counting. During inference, only the learnt embeddings are used, in a zero-shot manner, without knowledge of exemplars. As videos are of variable length and repetitions vary in duration, these approaches are not directly applicable to counting in videos.
Figure 1: VRC with ESCounts involves exemplars for relating information of the repeating action across the video. We visualise the density map with high relevance regions to the action push-up being highlighted, whilst regions of low relevance are not.
Taking inspiration from image-based approaches, we address Video Repetition Counting (VRC) using exemplars, for the first time. We differ from prior works that formulate VRC as classifying a preset number of repetitions [6, 16, 67], or consecutively detecting relevant parts of the repetition [15, 20, 30], such as the start/end. Instead, we argue that learning correspondences to reference exemplar(s) can provide a strong prior for discovering repetitions in videos. We propose Every-Shot Counts (ESCounts), a transformer-based encoder-decoder that encodes videos of varying lengths alongside exemplars and learns representations of general repeating motions, as shown in Figure 1. Similar to prior works [20, 30], we use the density map to regress the temporal location of each repetition. During training, we learn an exemplar latent representation, which we use for inference where no exemplars are available.
In summary, our contributions are as follows: (i) We introduce exemplar-based counting for VRC (ii) We propose an attention-based encoder-decoder that learns to correspond exemplars to a query video of varying length, and predict repetitions. (iii) We use a learnt latent for general motions in repetitions enabling our approach to predict the number of repetitions during inference without exemplars, (iv) We evaluate our approach on the three datasets commonly used for repetition counting in videos: RepNet [20], Countix [16], and UCFRep [65]. Our approach achieves a new state-of-the-art in every benchmark, even on Countix where the start-end times of repetitions are not annotated.
2 Related Works
We first review methods for the long-established task of object counting in images. We then review VRC methods for videos.
2.1 Object Counting in Images
Methods can be divided into class-specific and class-agnostic object counting.
Class-Specific Counting. These methods learn to count objects of singular classes or sets of categories e.g. people [56, 29], cars [19], wildlife [4]. A large portion of object-counting approaches [19, 56, 12, 40, 47] have relied on detecting target objects and counting their instances. Traditional methods have used hand-crafted feature descriptors to detect human heads [56] or head-shoulders [29] for crowd-counting. Other methods have used blobs [26], individual points [34], and object masks [13] for detecting and counting instances. Though object detection can be a preliminary step before counting, detection methods rely strongly on the performance of the object detector which can be less effective in densely crowded images [12]. A different set of methods has instead relied on regression, to either regress to the target count [12, 61] or estimate a density map [27, 41, 66].
Class-Agnostic Counting. Class-specific counting approaches are impractical for general setting, as prior knowledge of the object category is not available. Recent works [10, 3, 31, 48] have used one (or a few) exemplars as references to estimate a density map for unknown target classes. Exemplars for object counting were first introduced by Lu et al. [38]. Building on the property of image self-similarity[38] proposed a convolutional matching network and cast counting as an image-matching problem, where exemplar patches from the same image are used to match against other patches within the image. Following up, Liu et al. [10] proposed a transformer-based feature interaction module. They used an encoder for the query image and a convolution-based encoder for the exemplar. The interaction module cross-attended information between the exemplar and image and used a convolutional decoder to regress the density map. Recent approaches have also fused text and visual embeddings [3], used contrastive learning across modalities [24], and generated exemplar prototypes using stable diffusion [62]. Inspired by these methods, we propose a fully-attention-based encoder-decoder that extends exemplar-based counting to VRC. Our approach is invariant to video lengths and can use both learnt or encoded exemplars.
2.2 Video Repetition Counting (VRC)
Compared to image-based counting, video repetition counting has been less explored. Early approaches have compressed motion into a one-dimensional signal and recovered the repetition structure from the signal’s period [2, 25, 37, 42]. The periodicity can then be counted by Fourier analysis [5, 7, 14, 43, 2, 25], peak detection [52], or wavelet analysis [46]. However, these methods are limited to uniformly periodic repetitions. For non-periodic repetitions, localisation frameworks [11, 36, 49, 33] have been adapted. Zhang et al. [65] proposed a context-aware scale-insensitive framework to count repetitions of varying scales and duration. Their method exhaustively searched for pairs of consecutive repetitions followed by a prediction refinement module. Recent methods [15, 16, 20] have also extended image self-similarity to the temporal dimension with temporal self-similarity matrices (TSM). TSM is constructed using pair-wise similarity of embeddings over temporal locations. RepNet [16] used a transformer-based period predictor. To count repetitions with varying speeds, Trans-RAC [20] modified TSM to use multi-scale sequence embeddings. For counting under poor lighting conditions, [67] used audio in tandem with video within a multi-modal framework. They selectively aggregated information from the two modalities using a reliability estimation module. Li et al. [30] also used multi-modal inputs with optical flow as an additional signal supporting RGB for detecting periodicity. Yao et al. [63] proposed a lightweight pose-based transformer model that used action-specific salient poses as anchors and matched them across the video. However, the requirement of salient pose labels for each action limits its generalisability to unseen repetitions.
The above methods do not utilise the correspondences discovered by exemplar repetitions. Thus, do not relate variations in the action’s performance, alongside other visual variations such as viewpoint and camera motion. To handle such cases, we propose using action exemplars as references for VRC. Exemplars have previously been used in videos for action recognition tasks [17, 23, 59, 58]. [58] used silhouette/pose exemplars for classifying action sequences into predefined action categories using a Bayes classifier. [59] converted training videos to visual vocabulary and used the most discriminative visual words as exemplars for detecting actions. These methods are limited to a predefined set of classes and will be ineffective for class-agnostic tasks such as repetition counting. To our knowledge, we are the first to use exemplars for repetition counting in videos.
3 Every Shot Counts (ESCounts) Model
In this section, we introduce our ESCounts model (overviewed in Figure 2). We formally define encoding variable length videos alongside our model’s output, in Section 3.1. We then introduce the attention-based decoder that corresponds the input video to exemplars in Section 3.2. Predictions over temporally shifted inputs are combined for a final prediction, detailed in Section 3.3.
Figure 2: ESCounts Model overview. Bottom: Video $\mathbf{v}$ is encoded by $\mathcal{E}$ over sliding temporal windows to spatiotemporal latents $\mathbf{z}_{v}\in\mathbb{R}^{M\times C}$. Top Left: Exemplars $\{\mathbf{e}_{s}\}$ are also encoded with $\mathcal{E}$. Top Right: Video $\mathbf{z}_{v}$ and exemplar $\mathbf{z}_{s}$ latents are cross-attended by decoder $\mathcal{D}$ over $L$ cross-attention blocks. The resulting $\mathbf{z}_{L}\in\mathbb{R}^{M\times C}$ are attended over $L^{\prime}$ window self-attention blocks and projected into density map $\tilde{\mathbf{d}}$. The decoder $\mathcal{D}$ is trained to regress the error between predicted $\tilde{\mathbf{d}}$ and ground truth $\mathbf{d}$ density maps. At inference, the count is obtained by summing $\tilde{\mathbf{d}}$.
3.1 Input Encoding and Output Prediction
We denote the full video as $\mathbf{v}$ of varying $\mathcal{T}$ length and a fixed $H\times W$ spatial resolution. Segment $\mathbf{e}_{s}$ containing a single instance of the repeating action we wish to count, is selected as an exemplar. Exemplars are defined based on provided [[[[start, end]]]] labels of every repetition in the video [^1] For datasets where the start/end times are not available, pseudo-labels are used instead by uniformly dividing the video given the ground truth count.. During training, we select one or more exemplar shots $\mathcal{S}\subseteq\{\mathbf{e}_{1},\dots,\mathbf{e}_{s}\}$. Each training sample is thus a combination of the query video and a set of exemplars $(\mathbf{v},\mathcal{S})$.
We tokenise and encode the video $\mathbf{v}$ from its original size $\mathcal{T}\times H\times W$ into spatiotemporal latents $\mathbf{z}_{v}$. Encoder $\mathcal{E}$ is applied over a fixed-size sliding window, to account for the video’s variable length. The encoded video is represented by $\mathbf{z}_{v}\in\mathbb{R}^{M\times C}$ of $M=\mathcal{T^{\prime}}H^{\prime}W^{\prime}$ spatiotemporal resolution with $C$ channels. We note that $M$ is not a fixed number, as it depends on the video’s length $\mathcal{T}$. We add sinusoidal positional encoding to account for the relative order of these spatiotemporal latents while accommodating the variable video length.
During training, we select exemplars $\mathcal{S}$ from either the same video or another video containing the same action label; e.g. given a video containing push-up actions, we can sample exemplars from other videos showcasing the same action within the training set. We define a probability $p$ of sampling the exemplar from a different video; i.e. $p=0$ implies exemplars are only sampled from the same video, whereas for $p=1$ exemplars are always sampled from another video [^2] We ablate $p$ in our experiments.. In training, exemplars are sampled randomly from the labelled repetitions of the video. We use $\mathcal{E}$ to encode latent representations from each exemplar $\mathbf{e}_{s}\in\mathcal{S}$. We use the same encoder $\mathcal{E}$ for encoding $\mathbf{v}$ and $\mathbf{e}_{s}$ to allow correspondence.
We construct the ground truth density map $\mathbf{d}$ from the labelled repetitions in the video as a 1-dimensional vector. To match the downsampled temporal resolution of our input video $\mathcal{T^{\prime}}$, we also temporally downsample the ground-truth labels. The density map takes low values ($\approx 0$) at temporal locations without repetitions in the video, and high within repetitions. We use a normal distribution $\mathcal{N}$ centred around each repetition with $(\mu_{i}=\frac{t_{s}+t_{e}}{2},\sigma)$, where $t_{e}$ and $t_{s}$ are the start and end times of each repetition $i$.
$\mathbf{d}_{t}=\sum_{i}\mathcal{N}(t;\mu_{i},\sigma)\quad\forall\;t\in\{1,\dots,\mathcal{T}^{\prime}\}$ (1)
Note that the sum of the density map $\mathbf{d}$ matches the ground truth count, i.e. $\sum\mathbf{d}=c$ where $c$ is the ground truth count for the video.
3.2 Latent Exemplar Correspondence
Given both the encoded video $\mathbf{z}_{v}=\mathcal{E}(\mathbf{v})$ and exemplars $\mathbf{z}_{s}=\mathcal{E}(\mathbf{e}_{s})\;\forall\;\mathbf{e}_{s}\in\mathcal{S}$ latents, we use an attention-based decoder $\mathcal{D}(\mathbf{z}_{v},\mathbf{z}_{s})$ to attend to every repetition in the video $v$ that matches the encoded exemplar. Decoder $\mathcal{D}$ takes the encoded video $\mathbf{z}_{v}$ as input and predicts the location of every repetition in the video. The decoder outputs a 1-dimensional predicted density map of length $\mathcal{T}^{\prime}$ corresponding to the occurrences of the repeating action given the exemplars.
Figure 3: Cross-Attention block. Video latents $\mathbf{z}_{v}$ are self-attended and then cross-attended with latents $\mathbf{z}_{s}$ from each exemplar $s\in\mathcal{S}$ and the learnt latent $\mathbf{z}_{0}$ with the same weights. The resulting representations are then averaged.
Cross-attention Blocks. We explore the similarity between representations of exemplars and the query video to predict the corresponding locations of repetitions that match the exemplar. Thus, inspired by [10], we use cross-attention to relate exemplar and video latents. We define $L$ cross-attention blocks. Each block initially self-attends $\text{SA}(\cdot)$ the video latents $\mathbf{z}_{l}\in\mathbb{R}^{M\times C}$ with multi-head self-attention. We note that for the first layer, $\mathbf{z}_{1}=\mathbf{z}_{v}$. Then we relate exemplar and video information by cross-attending $\text{CA}(\cdot)$ video and exemplar latents. The block’s initial self-attention operation is formulated as:
$\mathbf{z}^{\prime}_{l}=\text{SA}(\text{LN}(\mathbf{z}_{l}))+\mathbf{z}_{l}\,\;\forall\,l\in\,\{1,\dots,L\},$ (2)
where $\text{LN}(\cdot)$ is layer normalisation.
[…]
$\mathbf{z}^{\prime\prime}_{l}=\frac{1}{N}\sum_{s=1}^{\mathcal{S}}\text{CA}(\mathbf{z}_{s},\text{LN}(\mathbf{z}^{\prime}_{l}))+\mathbf{z}^{\prime}_{l}\,\;\forall\,l\in\,\{1,\dots,L\},$ (3)
where $\mathcal{S}$ is the set of exemplars selected and $\mathbf{z}_{s}$ is the latent for the $s^{\text{th}}$ exemplar.
We also want to learn representations for repeating motions to estimate repetitions without explicitly providing exemplars. We thus define a learnable latent $\mathbf{z}_{0}$ to cross-attend $\mathbf{z}_{v}$. At each training iteration, we select exemplars from $\{\mathbf{z}_{0},\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$ and perform CA with $\mathbf{z}_{0}$ or $\{\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$. At inference, we use $\mathbf{z}_{0}$.
We use a multi-layer perceptron MLP on the exemplar-fused latents $\mathbf{z}^{\prime\prime}_{l}$.
$\mathbf{z}_{l+1}=\text{MLP}(\text{LN}(\mathbf{z}^{\prime\prime}_{l}))+\mathbf{z}^{\prime\prime}_{l}\,\;\forall\,l\in\,\{1,\dots,L\}$ (4)
The cross-attention blocks’ output is defined as $\mathbf{z}_{L}\in\mathbb{R}^{M\times C}$.
Window Self-attention Blocks. We explore the spatio-temporal inductive bias within the self-attention blocks. Each latent thus attends locally, to its spatio-temporal neighbouring tokens, over $L^{\prime}$ window self-attention $\text{WSA}(\cdot)$ [35] layers. We denote $\forall\,l\in\,\{L+1,\dots,L+L^{\prime}\}$ :
$\mathbf{z}_{l+1}=\text{MLP}(\text{LN}(\mathbf{z}^{\prime}_{l}))+\mathbf{z}^{\prime}_{l},\,\text{where}\;\mathbf{z}^{\prime}_{l}=\begin{cases}\text{WSA}(\text{LN}(\mathbf{z}_{l}))+\mathbf{z}_{l},\;\text{if}\;l=L+1\\ \text{WSA}(\text{shift}(\text{LN}(\mathbf{z}_{l})))+\mathbf{z}_{l},\;\text{else}\end{cases}$ (5)
where WSA is window self-attention. Note that following [35] windows are shifted at each layer to account for connections across different windows.
The output of the window self-attention blocks is of size $\mathbf{z}_{L+L^{\prime}}\in\mathbb{R}^{M\times C}$ In turn, $\mathbf{z}_{L+L^{\prime}}$ encodes repetition-relevant features over space and time. Thus, we use $\mathbf{z}_{L+L^{\prime}}$ to predict density map $\tilde{\mathbf{d}}$ for the occurrences of the target repeating action over time. We use a fully connected layer to project the latent to a 1-channel vector, i.e. MLP: ${\mathbb{R}^{M\times C}\rightarrow\mathbb{R}^{M}}$. We then pool the spatial resolution $H^{\prime}W^{\prime}$ resulting to predicted density map $\tilde{\mathbf{d}}\in\mathbb{R}^{\mathcal{T}^{\prime}}$.
Training Objective. Given ground-truth $\mathbf{d}$ and the predicted $\tilde{\mathbf{d}}=\mathcal{D}(\mathbf{z}_{v},\mathbf{z}_{s})$ density maps, we train $\mathcal{D}$ to regress the Mean Square Error between $\mathbf{d}$ and $\tilde{\mathbf{d}}$, and following [67], the Mean Absolute Error between ground truth counts $c$ and the predicted counts $\tilde{c}$ obtained by linearly summing the density map $\tilde{c}=\sum\tilde{\mathbf{d}}$
$\mathcal{L}=\underbrace{\frac{||\mathbf{d}-\tilde{\mathbf{d}}||^{2}}{\mathcal{T}^{\prime}}}_{\text{MSE}(\mathbf{d},\tilde{\mathbf{d}})}+\underbrace{\frac{|c-\sum{\tilde{\mathbf{d}}}|}{c}}_{\text{MAE}(c,\tilde{c})}$ (6)
Figure 4: Shifted Density maps from each video, are meaned to $\tilde{\mathbf{d}}$.
At inference, we use the predicted count $\tilde{c}$.
3.3 Time-Shift Augmentations
The predicted density map $\tilde{\mathbf{d}}$ results from encoding a video with $\mathcal{E}$, over non-overlapping sliding windows. However, as each window is of fixed temporal resolution, repetitions may span over multiple windows. Thus, we include time-shifting augmentations in which the start time of the encoded video is adjusted to allow for different spatiotemporal tokens. We train with augmentations of the start time whilst at inference, we use an ensemble of time-shift augmentations for a more robust estimation. We use multiple overlapping sequences as shown in Figure 4 and combine the predicted density maps over $\mathbf{K}$ shifted start/end positions. We obtain the final predicted density map by temporally aligning and averaging the predictions.
$\tilde{\mathbf{d}}_{t}=\frac{1}{|\mathbf{K}|}\underset{k\in\mathbf{K}}{\sum}\tilde{\mathbf{d}}^{k}_{t+\epsilon_{k}}$ (7)
where $\epsilon_{k}$ is the shifting for each $k\in\mathbf{K}$.
4 Experiments
We overview the used datasets, implementation details, and evaluation metrics in Sec. 4.1. We include quantitative and qualitative comparisons to state-of-the-art methods in Sec. 4.2. We ablate over different ESCounts settings in Sec. 4.3. We demonstrate ESCounts’ ability to utilise exemplars at inference, in Sec. 4.4.
4.1 Experimental Setup
Datasets. We evaluate our method on a diverse set of VRC datasets.
RepCount [20] contains videos of workout activities of varying repetition durations and appearances. Annotations include count labels alongside start and end timestamps per repetition. We use the publicly available Part-A of the dataset with $758$, $131$, and $152$ videos for train, val, and test respectively. We tune the hyperparameters on the val set and report our results on the test set.
Countix [16] is a subset of Kinetics [9] containing videos of repetitive actions and their associated counts. It is the largest VRC dataset with $4588$, $1450$, and $2719$ videos for train, val, and test respectively. Count annotations are provided per video, without individual repetition start-end times. We use pseudo-repetition annotations by dividing the videos into uniform segments based on the ground truth count, as Countix does not have many pauses or interruptions between counts. We use the pseudo-labels for estimating the density maps, without additional annotations, to compare directly to other methods.
UCFRep [65] is a subset of UCF-101 [51] consisting of 420 videos for train and 106 videos for val from 23 categories. Following [65, 30], we report our results on the val split as no test set is available. Like RepCount, it includes annotations of start and end timestamps for each repetition.
Implementation Details. Unless specified otherwise, we use VideoMAE-VIT-B [55] as our encoder $\mathcal{E}$ with weights pre-trained on Kinetics-400 [9]. We sample frames from variable-length videos every 4 frames using a sliding window of $64$ frames. At each window, our encoder’s input is of $16\times 224\times 224$ size, and the output is $8\times 14\times 14$, resulting in $1568$ spatiotemporal tokens. We use $C=512$ channels333when encoders have a different output, we add a fully connected layer to map to $C$. The input to the decoder is of variable length $M=1568\frac{\mathcal{R}}{64}$, where $\mathcal{R}$ is the total number of frames in the video at raw framerate. Exemplars are sampled uniformly with $16$ frames between the start and end of a repetition.
The encoder is frozen with only the decoder and the zero-shot latent $\mathbf{z}_{0}$ being trained. We use $L=2$ and $L^{\prime}=3$. We ablate $L$ and $L^{\prime}$ in the supplementary material. We train our model on a single Tesla V100 for 300 epochs with a batch size of 1, to deal with variable-length videos, and accumulate gradients over 8 batches. We use an initial learning rate of $55e-$ 5 and decay it every $60$ epochs by $0.8$, and a weight decay of $25e-$ 2. For each training instance, we randomly set the number of exemplars $|\mathcal{S}|\sim\{0,1,2\}$ and sample $\mathcal{S}$ exemplars. We set the chance of sampling exemplars from a different video to $p=0.4$.
At inference, we only use the learnt latent to predict repetition counts. We aggregate predictions over $|\mathbf{K}|=4$ sequences.
Evaluation Metrics. Following previous repetition counting works [20, 16, 67], we use Mean Absolute Error (MAE) and Off-By-One accuracy (OBO) as evaluation metrics, calculated as Eqs. 8 and 9 respectively. Inspired by image counting methods [10, 3], we introduce Root-Mean-Square-Error (RMSE) in Eq. 10 for VRC providing a more robust metric for diverse counts compared to MAE’s bias towards small counts. With continuous improvements in repetition counting methods, we also report the off-by-zero accuracy (OBZ) in Eq. 11 as a tighter metric than the corresponding OBO for precise counts.
$MAE=\frac{1}{|\Omega|}\sum_{i\in\Omega}\frac{|c_{i}-\tilde{c}_{i}|}{c_{i}},$ (8)
$OBO=\frac{1}{|\Omega|}\sum_{i\in\Omega}\mathds{1}(|c_{i}-\tilde{c}_{i}|\leq 1),$ (9)
$RMSE=\sqrt{\frac{1}{|\Omega|}\sum_{i\in\Omega}(c_{i}-\tilde{c}_{i})^{2}},$ (10)
$OBZ=\frac{1}{|\Omega|}\sum_{i\in\Omega}\mathds{1}(|c_{i}-\tilde{c}_{i}|=0),$ (11)
where $c_{i}$, $\tilde{c}_{i}$ are the ground-truth and predicted counts for $i$-th video in test set $\Omega$. $\mathds{1}$ is the indicator function.
4.2 Comparison with State-of-the-art
In Tab. 2, we compare our method, ESCounts, to prior methods on the three datasets. We also provide results on the same backbone as the best-performing method on each dataset, for fair and direct comparison to previous works.
Table 1: Comparison of VRC methods. $\dagger$ represents multi-modal models that use added audio or flow. denotes results reproduced using provided checkpoints. denotes inhouse retraining using published codes. Grayed rows in (c) represent methods that finetune the encoder. Top performances for each metric and dataset are in bold.
(a) RepCount
[…]
RepNet [16] R2D50 - 0.995 - 0.013
TransRAC [20] VSwinT  9.130 0.443  0.085 0.291
MFL [30$\dagger$ VSwinT - 0.384 - 0.386
ESCounts VSwinT 6.905 0.298  0.1830.183\;0.183\;0.183 0.403
ESCounts VMAE  $\;\mathbf{4.455}\;$ $\mathbf{0.213}$  $\;\mathbf{0.245}\;$ $\mathbf{0.563}$
(b) Countix
[…]
RepNet[16] R2D50 - 0.364 - 0.697
Sight & Sound [67$\dagger$ R(2+1)D18 - 0.307 - 0.511
ESCounts R(2+1)D18 3.536 0.293 0.286 $\mathbf{0.701}$
ESCounts VMAE $\mathbf{3.029}$ $\mathbf{0.276}$ $\mathbf{0.319}$ 0.673
(c) UCFRep
[…]
Levy & Wolf [28] RX3D101 - 0.286 - 0.680
RepNet [16] R2D50 - 0.998 - 0.009
Context (F) [65] RX3D101 5.761 0.653 0.143 0.372
TransRAC [20] VSwinT - 0.640 - 0.324
MFL[30$\dagger$ RX3D101 - 0.388 - 0.510
ESCounts RX3D101 2.004 0.247 0.343 0.731
ESCounts VMAE $\mathbf{1.972}$ 0.216 0.381 0.704
Context [65] RX3D101 2.165 0.147 $\mathbf{0.452}$ 0.790
Sight & Sound [67$\dagger$ R(2+1)D18 - $\mathbf{0.143}$ $\mathbf{0.800}$
Table 2: Cross-dataset generalisation scores. Arrows $X\rightarrow Y$ denote train dataset $X$ and test dataset $Y$. Results obtained using provided checkpoints are denoted with .
RepCount $\rightarrow$ UCFRep RepCount $\rightarrow$ Countix
[…]
RN [16] - 0.998 - 0.009 - - - -
TRAC [20] 6.701 0.640 0.087 0.324 6.867 0.593 0.132 0.364
MFL [30] - 0.523 - 0.350 - - - -
ESCounts $\mathbf{3.536}$ $\mathbf{0.317}$ $\mathbf{0.219}$ $\mathbf{0.571}$ $\mathbf{4.429}$ $\mathbf{0.374}$ $\mathbf{0.185}$ $\mathbf{0.521}$
Tab. 2 - RepCount. We demonstrate that our ESCounts model outperforms recent methods [20, 30] with the same Swin Tiny (VSwinT) [35] encoder, by $-0.15$ in MAE and $+0.11$ in OBO. Our approach also improves upon [30] which uses optical flow and video in tandem, by margins of $-0.09$ in MAE and $+0.02$ in OBO. The largest improvements are observed with VideoMAE (VMAE) surpassing [20] on the introduced OBZ by $+0.16$ and reducing RMSE by $-4.68$.
Tab. 2 - Countix. Compared to the state-of-the-art [16, 67] our ESCounts consistently outperforms other models with the same R(2+1)D18 encoder. Our video-only model surpasses the audio-visual model in [67] by $+0.19$ OBO. Further improvements on the RMSE, MAE, and OBZ are observed with VMAE.
Tab. 2 - UCFRep. Across approaches with frozen encoders ESCounts improve the top-performing [30] by $+0.19$ OBO and $-0.17$ MAE. Our method does not outperform [67, 65] that fine-tune their encoders on the training set of UCFRep which is advantageous given the dataset’s size as also noted in [30]. We showcase this with Context (F) trained from the publicly available code of [65] with a frozen encoder and result in a significant drop in performance with +0.51 MAE and -0.42 OBO. Despite this, we achieve strong results.
Figure 5: Scatter plot, instances, and density maps from RepCount, Countix and UCFRep. The dotted diagonal denotes correct predictions. We compare ESCounts against TransRAC [20] on Repcount and Context [65] on UCFRep. Examples a--m are marked with corresponding action class and count predictions. We add the Ground Truth (GT) and Predicted (P) density maps for each instance. For Countix, pseudo-labels are shown as GT, as no start-end times are given.
Qualitative Results. In Fig. 5 we visualise predicted to ground truth counts. For RepCount and UCFRep, we select [20] and [65] as respective baselines and use their publicly available checkpoints444[30] was not used as a baseline as the code is not public. The publicly available checkpoints on Countix obtained lower results than originally reported.. ESCounts accurately predicts the number of repetitions for a wide range of counts, with most predictions being close to the diagonal. Though predictions from both the baseline and ESCounts are close to the ground truth in low counts, they significantly diverge in high counts. We additionally visualise specific examples with sample frames from these videos. ESCounts is robust to the magnitude of counts, with accurate predictions over low count (a,b,g,k,l) and high count (d,i,m) examples. In cases of over- and under-predictions; e.g. (c,e,f,h,j,n) ESCounts predictions remain closer to actual counts. As shown by the density maps, ESCounts can also localise the repetitions. For Countix, because the model was trained on pseudo labels, even though ESCounts can predict accurate counts, it struggles to localise the repetitions at times. We further investigate the localisation capabilities in the supplementary.
Cross-dataset Generalisation. Following prior works [20, 30], we also test the generalisation capabilities of our method. We use our model trained on RepCount and evaluate it on the test sets of UCFRep and Countix. As shown in Tab. 2, our method achieves the best results surpassing [30] by $-0.21$ in MAE and $+0.22$ in OBO in UCFRep. We also explore cross-dataset generalisation for Countix, and select [20] as a baseline with the public checkpoint. We outperform [20] by significant margins across metrics, showcasing the strong ability of ESCounts to generalise across datasets.
4.3 Ablation Studies
In this section, we conduct ablation studies on RepCount [20] using VMAE as the encoder. We study the impact of exemplars by replacing cross- with self-attention and vary the number of exemplars used during training. We evaluate the sensitivity of our method to the exemplar sampling probability $p$. We ablate $\sigma$ in density map estimation and analyse the impact of time-shift augmentations and our loss function’s components.
Table 3: Ablations on RepCount over different ESCounts settings.
(a) SA-only decoder
[…]
SA-only 5.654 0.273 0.147 0.470
ESCounts $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
(b) Number of exemplars $|\mathcal{S}|$
[…]
$|\mathcal{S}|=0$ 4.962 0.240 0.223 0.519
$|\mathcal{S}|\sim\{0,1\}$ 4.633 0.228 0.236 0.546
$|\mathcal{S}|\sim\{0,2\}$ 4.601 0.226 0.239 0.550
$|\mathcal{S}|\sim\{0,1,2\}$ $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
(c) Exemplar sampling
Diff same
[…]
- 4.701 0.224 0.226 0.521
5.553 0.270 0.165 0.464
$\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
(d) Sampling prob. $p$
[…]
$0.0$ 4.919 0.240 0.205 0.545
$0.2$ 4.654 0.221 0.236 0.550
$0.4$ $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
$0.6$ 4.561 0.218 0.240 0.558
$0.8$ 4.735 0.230 0.223 0.553
$1.0$ 5.012 0.245 0.218 0.532
(e) Density peaks $\sigma$
[…]
Variable 6.152 0.301 0.165 0.457
0 5.145 0.241 0.206 0.510
0.25 4.871 0.226 0.228 0.542
0.50 $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
0.75 4.683 0.218 0.240 0.556
1.00 4.732 0.223 0.238 0.552
(f) Timeshift Aug. $|\mathbf{K}|$
[…]
1 4.592 0.221 0.235 0.552
2 4.493 0.217 0.242 0.556
3 4.471 0.213 0.243 0.561
4 $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
(g) Effect of Objective
[…]
MSE 5.109 0.273 0.215 0.532
[…]
Do exemplars help in training? We study the impact of using exemplars for training by directly replacing the cross-attention decoder blocks with self-attention. As seen in Tab. 3, using self-attention (SA-only) performs significantly worse than our proposed ESCounts. Cross-attending exemplars decrease the RMSE/MAE by $-1.20$ and $-0.06$ whilst improving OBZ and OBO by $+0.10$ and $+0.09$, thus emphasising the benefits of exemplar-based VRC.
How many exemplars to sample? Since we use a varying number of exemplars whilst training, we also experiment by varying $|\mathcal{S}|$ from $0-2$ in Tab. 3. In $|\mathcal{S}|=0$ we train only with the learnt zero-shot latent. Training with $|\mathcal{S}|\sim\{0,1,2\}$ provides the best zero-shot scores in inference. We believe training with a larger and more diverse set of exemplars helps the model to learn a general zero-shot token that can capture general information about repetitions. Due to memory constraints, we could not train with more than 2 exemplars.
How to sample exemplars? We analyse the impact of sampling exemplars from other videos, from the same or a different action class. As expected, Tab. 3 shows that keeping the same action category for both exemplars and query video performs the best, as ensuring the same action semantics between exemplars and query video helps to learn their correspondence. In this table, we used sampling probability $p=0.4$.
We vary the sampling probability from other videos of the same underlying action $p$ in Tab. 3. For $p=0.0$ exemplars are sampled exclusively from the query video, whilst for $p=1.0$ exemplars are sampled solely from other videos of the same class. The best performance was observed with $p=0.4$, showcasing sampling from the same video to be critical for learning.
[…]
OBZ
Ground truth number of repetitionsRepetition lengthVideo duration
(a) Off by N (OBN)
(b) Class-wise OBZ
(c) RMSE
(d) MAE
(e) OBZ
(f) OBO
(g) RMSE
(h) MAE
(i) OBZ
(j) OBO
(k) RMSE
(l) MAE
(m) OBZ
(n) OBO
Figure 6: Grouped VRC scores over different number of repetitions and lengths. (a) overviews the Off by N accuracy for increasing Ns. (b) shows OBZ by action class. The first row (c–f) reports results over different counts. (g–j) reports scores over groups by repetition durations. (k–n) reports metrics grouped by video duration.
What’s the impact of time-shift augmentations? As predictions are aggregated over $|\mathbf{K}|$ density maps by time-shifting the video input. As shown in Tab. 3, having $|\mathbf{K}|=4$ shifted start/end positions provides the best results. However, results are strong even without test-time augmentations in $|\mathbf{K}|=1$.
What should be the density map variance? Density maps are constructed as vectors with normal distributions $\mathcal{N}(\cdot;\mu,\sigma)$ over repetition starts/ends timestamps. Reducing $\sigma$ increases the sharpness, resulting in a single delta function for $\sigma=0$. We ablate over different $\sigma$ in Tab. 3. Denser and successive repetitions can benefit from sharper peaks of small $\sigma$ and sparser repetitions of larger durations can benefit from large $\sigma$. We also ablate using variable $\sigma$ that changes with the duration of repetition segments. Having $\sigma=0.5$ provides the best results with a balance between sharpness and covering the duration of repetitions.
How helpful is the MAE for the objective? We analyse ESCounts’ performance with and without the MAE loss from [67] in Tab. 3. The combined objective helps performance for diverse counts across all metrics.
How close are predictions to the ground truth?. We further relax the off-by metrics to Off-By-N in Figure 6(a) to visualise the proximity of predictions to the ground truth. Overall, 84% of predictions are within ±$\pm 3$ of the actual count.
What is the performance per action category?. In Figure 6(b), we plot the OBZ per action class. ESCounts performs fairly uniformly across all classes with the best performing categories being pommelhourse and squat.
How does performance differ across counts, repetition lengths, and video durations? Up to this point, we have focused on the performance across all videos regardless of individual attributes. We now consider the sensitivity of ESCounts across equally sized groups based on the number of repetitions, average repetition length, and video duration.
We report all metrics over groups of counts in Figs. 6(c), 6(d), 6(e) and 6(f). As expected, our method performs best in groups of smaller counts with higher counts being more challenging to predict precisely.
In Figs. 6(g), 6(h), 6(i) and 6(j) we report VRC metrics with results grouped by average video repetition duration. These are grouped, into equal sized bins, to XS=(0-0.96)s, S=(0.96-1.53)s, M=(1.53-2.29)s, L=(2.29-3.09)s, XL=(>3.09)s. Predicting density maps is more challenging for short repetitions. However, ESCounts can still correctly predict counts across repetition lengths as shown by Figures 6(i) and 6(j). We also group videos by duration into XS=(8.0-11.0)s, S=(11.0, 26.0)s, M=(26.0, 33.9)s, L=(33.9-45.9)s and XL=(45.9-68.0)s. From Figures 6(k), 6(l), 6(m) and 6(n), counting repetitions from longer videos is more challenging.
Table 4: Number of shots at inference. We test using exemplars from the same video or a different video of the same action class from the train set.
RepCount UCFRep
[…]
0 N/A 4.455 0.213 0.245 0.563 1.972 0.216 0.381 0.704
1 4.432 0.207 0.251 0.563 1.912 0.211 0.388 0.712
4.369 0.210 0.247 0.589 1.890 0.203 0.400 0.714
2 4.384 $\mathbf{0.206}$ 0.251 0.572 1.885 0.208 0.391 0.720
4.360 0.209 0.247 0.592 1.857 0.199 0.419 0.718
3 4.381 0.207 $\mathbf{0.252}$ 0.579 1.878 0.207 0.399 $\mathbf{0.730}$
$\mathbf{4.351}$ $\mathbf{0.206}$ 0.250 $\mathbf{0.596}$ $\mathbf{1.855}$ $\mathbf{0.198}$ $\mathbf{0.420}$ 0.723
4.4 Inference with More Shots
up to this point, only the learnt zero-shot latent was used for inference.
Prior works in object counting [10, 38] have reported results using exemplars (i.e. object crops) in inference. While this is not comparable to other works in VRC, we can assess our method’s ability to utilise exemplars during inference in Tab. 4. Video exemplars steadily improve performance as the number of exemplars increases. As our model cross-attends exemplars in parallel, we can train with $0-2$ exemplars, and use $>2$ exemplars during inference. We show comparable results when sampling exemplars from the test video vs exemplars from the training set videos showcasing the same action category. Combined with a classifier, one can envisage an approach that classifies the action then sources exemplars from the training set to assist counting during inference.
5 Conclusion
We have proposed to utilise exemplars for video repetition counting. We introduce Every Shot Counts (ESCounts), an attention-based encoder-decoder that learns to correspond exemplar repetitions across a full video. We define a learnable zero-shot latent that learns representations of generic repetitions, to use during inference. Extensive evaluation on RepCount, Countix, and UCFRep demonstrates the merits of ESCounts achieving state-of-the-art results on the traditional MAE and OBO metrics and the newly introduced RMSE and OBZ. Further analysis showed that the exemplars do not need to be from the same video as long as they belong to the same action class. The diversity of these exemplars is an aspect for future exploration.
Acknowledgements
This work uses publicly available datasets and annotations to produce all results and ablations. Research is supported by EPSRC UMPIRE (EP/T004991/1). S. Sinha is supported by EPSRC DTP (Doctoral Training Program) studentship.
Appendix
Code is made available anonymously from: https://github.com/sinhasaptarshi/EScounts. We provide the full code, checkpoint as well as a demo with a couple of videos to try our inference.
In this supplementary, we explain the attached qualitative video in Sec. 6. We then provide additional ablations on the architecture’s choices (e.g. depth of transformer and window size) in Sec. 7. Additionally, we evaluate the ability of ESCounts to locate each repetition within the video in Sec. 8. Finally, we analyse our improvement over baseline by studying groups of videos based on their length, number of repetitions, and counts in Sec. 9.
6 Qualitative Video and Extended Figure
We attach a compilation video (1380.mp4) showcasing our method’s abilities in Video Repetition Counting (VRC). For a diverse set of 11 videos from all 3 datasets, we show the video alongside synchronised ground truth and ESCounts predicted density maps. For each example, we note which test set it is from.
Additionally, we extend Fig. 5 in the main paper with additional examples in Fig. 8.
7 Further Ablations
We extend the ablation studies in  Sec. 4.3 by reporting results over different $L$ and $L^{\prime}$. Additionally, we analyse the impact of windowed-self attention on the performance of ESCounts.
(o) RepCount
(p) Countix
Figure 7: Additional qualitative results.
(a) UCFRep
Figure 8: Additional qualitative results (continued).
Table 5: Impact of $L$.
[…]
1 4.843 0.229 0.223 0.545
2 $\mathbf{4.455}$ $\mathbf{0.213}$ 0.245 $\mathbf{0.563}$
3 4.575 0.219 $\mathbf{0.247}$ 0.560
4 4.783 0.225 0.235 0.548
Table 6: Impact of $L^{\prime}$.
[…]
1 4.932 0.247 0.212 0.525
2 4.634 0.218 0.238 0.550
3 $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
4 4.532 0.225 0.230 0.552
Table 7: Window sizes.
[…]
$(3,3,3)$ 5.212 0.261 0.185 0.521
$(2,7,7)$ 4.871 0.247 0.201 0.537
$(4,7,7)$ $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
$(7,7,7)$ 4.753 0.225 0.232 0.520
full 5.011 0.227 0.221 0.533
Impact of $L$. We ablate the value of $L$ i.e. the number of layers in the cross-attention block. Increasing $L$ increases the number of operations that discover correspondences between the video and the selected exemplars. As seen in Tab. 7, while low $L$ causes a drop in performance, high $L$ can also be detrimental probably due to overfitting. $L=2$ gives the best results for the majority of the metrics.
Next keeping $L=2$ fixed, we vary $L^{\prime}$ in Tab. 7. $L^{\prime}$ is the number of windowed self-attention layers in the self-attention block. $L^{\prime}=3$ gives the best results on all the metrics. Similarly, increasing or decreasing $L^{\prime}$ drops performance gradually.
Self-attention vs Windowed Self-attention. Motivated by [35], we use windowed self-attention for the self-attention block in ESCounts. Given spatio-temporal tokens $\mathcal{T}^{\prime}\times H^{\prime}\times W^{\prime}\times C$, windowed self-attention computes multi-headed attention for each token within the immediate neighbourhood using 3D shifted windows of size $t^{\prime}\times h^{\prime}\times w^{\prime}$, where $t^{\prime}<\mathcal{T}^{\prime}$, $h^{\prime}<H^{\prime}$ and $w^{\prime}<W^{\prime}$. Here we ablate for various values $(t^{\prime},h^{\prime},w^{\prime})$ in Tab. 7. Note that $t^{\prime}=\mathcal{T}^{\prime}$, $h^{\prime}=H^{\prime}$ and $w^{\prime}=W^{\prime}$ is using standard self-attention over all the embeddings, where each token is attended with every token and is denoted by full. As shown, the best performance is obtained with window size $(4,7,7)$, demonstrating the importance of attending to tokens in immediate spatio-temporal neighbourhoods only. We found that the results are most sensitive to the value of $t^{\prime}$. Performance steadily drops as the value of $t^{\prime}$ changes.
Sampling Rate for Encoding. As stated in the implementation details, we construct our input by sampling every four frames from the video. We ablate the impact of the sampling rate in Tab. 8. From the results, we observe that denser sampling is key for robust video repetition counting. Reducing the sampling rate steadily decreases performance.
Table 8: Impact of sampling rate
Sampling every
$n$ frames
[…]
4 4.455 0.213 0.245 0.563
8 5.112 0.268 0.221 0.521
16 5.911 0.296 0.185 0.482
32 6.562 0.346 0.156 0.444
Table 9: OBO, parameters, and training and inference speeds on UCFRep. Mtreics obtained by the public available codebase of [65] are denoted with .
[…]
#Trainable
params (M)
Train set $\downarrow$
(sec/sample)
Test set $\downarrow$
(sec/sample)
Context[65] 0.790 47.6 1.171 1.818
ESCounts 0.731 21.1 (-26.5) 0.138 (-1.033) 0.141 (-1.677)
Model Size and Speed For UCFRep [65], [65, 67] achieve better performance than ESCounts. However, this performance is achieved by having more trainable parameters, as [65, 67] finetune the encoders on the target dataset. We use the provided codebase from [65] and benchmark the average number of iterations per second for a full forward and backward pass over the entire training set alongside inference-only average times on the test set. We use the same experiment set-up described in Sec. 4.1 and report speeds in Table 9. Training ESCounts is  8 times faster. Interestingly, even during inference ESCounts is  12 times faster than Context [65] which uses iterative processing. Note that [67] could not be used for this analysis as their code for training with UCFRep is not publicly available.
[…]
Method $\theta$ values for relative threshold $r$
0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 Avg
Baseline [20] 38.59 37.46 35.02 32.55 30.40 26.97 22.66 17.22 12.17 28.12
ESCounts 38.83 38.64 38.07 37.44 35.82 33.43 30.76 27.52 20.85 33.48
8 Repetition Localisation
Figure 9: Explaining localisation metric $\mathcal{J}$. We identify local maxima in $\tilde{\mathbf{d}}$ and threshold peaks higher than $r$ to remove noise. $\mathcal{J}$ is then computed between the annotated start-end times and the thresholded peaks.
VRC metrics only relate predicted to correct counts, regardless of whether the repetitions have been correctly identified. We thus investigate whether the peaks of the predicted density map $\tilde{\mathbf{d}}$ align with the annotated start-end times of repetitions in the ground truth. Following action localisation methods [8, 18, 22], we adopt and adjust the Jaccard index $\mathcal{J}$ for repetition localisation. Using the predicted density map we define different thresholds for salient regions of repetitions. As the values of $\tilde{\mathbf{d}}$ peaks vary across videos, we apply thresholds $\theta$ relative to the maximum and minimum values, $r=\theta(\max(\tilde{\mathbf{d}})-\min(\tilde{\mathbf{d}}))$. We find all local maxima in $\tilde{\mathbf{d}}$ and only keep those above threshold $r$. We consider a repetition to be correctly located (TP) if at least one peak occurs within the start-end time of that repetition. Peaks that occur within the same repetition are counted as one. In contrast, peaks that do not overlap with repetitions are false positives (FP) and repetitions that do not overlap with any peak are false negatives (FN). We then calculate the Jaccard index as TP divided by all the correspondences (TP + FP + FN) as customary.
(a) counts
(b) repetition length
(c) video duration
Figure 10: Comparison of ESCounts and TransRAC [20] for groups of counts/repetition duration/video-length
In Table 10 we report the Jaccard index over different thresholds alongside the Mean Average Precision (mAP) on RepCount. We select TransRAC [20] as a baseline due to their publicly available checkpoint. Across thresholds, ESCounts outperforms [20] with the most notable improvements observed over higher threshold values demonstrating that ESCounts predict density maps with higher contrast between higher and lower salient regions. For 0.9, 0.8, and 0.7 thresholds ESCounts demonstrates an +8.68%, +10.30%, and +8.10% improvement over [20].
9 Comparison with TransRAC over Groups
As in Sec. 4.3, we divide the videos into groups based on their counts, average repetition duration, and lengths. We compare ESCounts’ performance against TransRAC in each of these groups. As evident from Fig. 10, ESCounts performs significantly better than TransRAC on high counts. From Fig. 10, ESCounts performs strongly over varying repetition duration, whilst TransRAC struggles on videos with XS and XL repetitions. Fig. 10 shows that even though both TransRAC and ESCounts performance drops on longer videos, ESCounts perform significantly better on them.

### v2

@@
 Extensive experiments over commonly used datasets (RepCount, Countix, and UCFRep) showcase ESCounts obtaining state-of-the-art performance across all three datasets.
-On RepCount, ESCounts increases the off-by-one from $0.39$ to $0.56$ and decreases the mean absolute error from $0.38$ to $0.21$.
 Detailed ablations further demonstrate the effectiveness of our method.
@@
 1 Introduction
-Recent years have come with tremendous progress in video understanding.
-Large Visual Language Models (VLMs) have been adopted for many vision tasks including video summarisation [44, 32, 60], localisation [45, 57], and question answering (VQA) [1, 21, 39, 64].
-Despite their great success, recent analysis [24] shows that VLMs can still fail to correctly count objects or actions.
-Robust counting can be challenging due to appearance diversity, limited training data, and most critically semantic ambiguity – identifying ‘what’ exactly to count.
-Evidence in developmental psychology and cognitive neuroscience [50, 53, 54] shows that infants fail to differentiate the number of hidden objects when they are not shown and counted to them first, suggesting an upper limit of individual objects that can be represented in working memory.
-However, infants that were first exposed to an instance of the object could better approximate the carnality.
-This shows that counting is a visual exercise of matching to given exemplars, and is developed in humans prior to understanding their semantics.
+In recent years, tremendous progress has been made in video understanding.
+Visual Language Models (VLMs) have been adopted for many vision tasks including video summarisation [49, 36, 65], localisation [50, 62], and question answering (VQA) [1, 23, 44, 69].
+Despite their great success, recent analysis [27] shows that VLMs can still fail to count objects or actions correctly.
+Robust counting can be challenging due to appearance diversity, limited training data, and the semantic ambiguity of identifying ‘what’ to count.
+Evidence in developmental psychology and cognitive neuroscience [56, 59, 60] shows that infants fail to differentiate the number of hidden objects if not shown and counted to them first, suggesting an upper limit of individual objects in working memory.
+However, infants exposed to an instance of the object first could better approximate carnality.
+This shows that counting is a visual exercise of matching to exemplars, and is developed before understanding their semantics.
 Object-counting in images has recently exploited exemplars to improve performance [43, 38].
 In training, models attend to one or more exemplars of ‘what’ object(s) to count alongside learnt embeddings for exemplar-free counting.
-During inference, only the learnt embeddings are used, in a zero-shot manner, without knowledge of exemplars.
-As videos are of variable length and repetitions vary in duration, these approaches are not directly applicable to counting in videos.
+During inference, only the learnt embeddings are used for zero-shot counting without knowledge from exemplars.
+As videos are of variable length and repetition durations vary, these approaches are not directly applicable to videos.
 Figure 1: VRC with ESCounts involves exemplars for relating information of the repeating action across the video.
 We visualise the density map with high relevance regions to the action push-up being highlighted, whilst regions of low relevance are not.
-Taking inspiration from image-based approaches, we address Video Repetition Counting (VRC) using exemplars, for the first time.
-We differ from prior works that formulate VRC as classifying a preset number of repetitions [6, 16, 67], or consecutively detecting relevant parts of the repetition [15, 20, 30], such as the start/end.
-Instead, we argue that learning correspondences to reference exemplar(s) can provide a strong prior for discovering repetitions in videos.
-We propose Every-Shot Counts (ESCounts), a transformer-based encoder-decoder that encodes videos of varying lengths alongside exemplars and learns representations of general repeating motions, as shown in Figure 1.
-Similar to prior works [20, 30], we use the density map to regress the temporal location of each repetition.
-During training, we learn an exemplar latent representation, which we use for inference where no exemplars are available.
-In summary, our contributions are as follows: (i) We introduce exemplar-based counting for VRC (ii) We propose an attention-based encoder-decoder that learns to correspond exemplars to a query video of varying length, and predict repetitions.
-(iii) We use a learnt latent for general motions in repetitions enabling our approach to predict the number of repetitions during inference without exemplars, (iv) We evaluate our approach on the three datasets commonly used for repetition counting in videos: RepNet [20], Countix [16], and UCFRep [65].
-Our approach achieves a new state-of-the-art in every benchmark, even on Countix where the start-end times of repetitions are not annotated.
+Taking inspiration from image-based approaches, we address Video Repetition Counting (VRC) with exemplars for the first time.
+We differ from prior works that formulate VRC as classifying a preset number of repetitions [6, 15, 72], or detecting relevant parts (start/end) of repetitions [14, 22, 33].
+Instead, we argue that learning correspondences to reference exemplar(s) during training can provide a strong prior for discovering correspondences across repetitions at inference.
+We propose Every-Shot Counts (ESCounts), a transformer-based encoder-decoder that during training encodes videos of varying lengths alongside exemplars and learns latents of general repeating motions, as shown in Fig.
+1.
+Similar to [22, 33], we use density maps to regress the temporal location of repetitions.
+At inference, learnt latents are used for exemplar-free counting.
+In summary, our contributions are as follows: (i) We introduce exemplar-based counting for VRC (ii) We propose an attention-based encoder-decoder that corresponds exemplars to a query video of varying length.
+(iii) We learn latents for general repeating features and use them to predict the number of repetitions during inference without exemplars, (iv) We evaluate our approach on the three commonly-used VRC datasets: RepNet [22], Countix [15], and UCFRep [70].
+Our approach achieves a new state-of-the-art in every benchmark, even on Countix where start-end times of repetitions are not annotated.
 2 Related Works
@@
 These methods learn to count objects of singular classes or sets of categories e.g.
-people [56, 29], cars [19], wildlife [4].
+people [61, 32], cars [21], or wildlife [4].
 A large portion of object-counting approaches [21, 61, 11, 45, 52] have relied on detecting target objects and counting their instances.
@@
 Other methods have used blobs [29], individual points [39], and object masks [12] for detecting and counting instances.
-Though object detection can be a preliminary step before counting, detection methods rely strongly on the performance of the object detector which can be less effective in densely crowded images [12].
-A different set of methods has instead relied on regression, to either regress to the target count [12, 61] or estimate a density map [27, 41, 66].
+Though object detection can be a preliminary step before counting, detection methods rely strongly on the object detector’s performance which can be less effective in densely crowded images [11].
+Other methods instead relied on regression, to either regress to the target count [11, 66] or estimate a density map [30, 46, 71].
 Class-agnostic counting.
-Class-specific counting approaches are impractical for general setting, as prior knowledge of the object category is not available.
+Class-specific counting approaches are impractical for general settings where prior knowledge of the object category is not available.
 Recent works [38, 3, 35, 54] have used one (or a few) exemplars as references to estimate a density map for unknown target classes.
-Exemplars for object counting were first introduced by Lu et al.
-[38].
-Building on the property of image self-similarity[38] proposed a convolutional matching network and cast counting as an image-matching problem, where exemplar patches from the same image are used to match against other patches within the image.
+Building on the property of image self-similarity[43] proposed a convolutional matching network.
+They cast counting as an image-matching problem, where exemplar patches from the same image are used to match against other patches within the image.
 Following up, Liu et al.
-[10] proposed a transformer-based feature interaction module.
-They used an encoder for the query image and a convolution-based encoder for the exemplar.
-The interaction module cross-attended information between the exemplar and image and used a convolutional decoder to regress the density map.
+[38] used an encoder for the query image, a convolution-based encoder for the exemplar, and an interaction module to cross-attend information between the exemplar and the image.
+A convolutional decoder was used to regress the density map.
 Recent approaches have also fused text and visual embeddings [3], used contrastive learning across modalities [27], and generated exemplar prototypes using stable diffusion [67].
-Inspired by these methods, we propose a fully-attention-based encoder-decoder that extends exemplar-based counting to VRC.
+Inspired by these methods, we propose an attention-based encoder-decoder that extends exemplar-based counting to VRC.
 Our approach is invariant to video lengths and can use both learnt or encoded exemplars.
@@
 Early approaches have compressed motion into a one-dimensional signal and recovered the repetition structure from the signal’s period [2, 28, 42, 47].
-The periodicity can then be counted by Fourier analysis [5, 7, 14, 43, 2, 25], peak detection [52], or wavelet analysis [46].
+The periodicity can then be counted with Fourier analysis [5, 7, 13, 48, 2, 28], peak detection [58], or wavelet analysis [51].
 However, these methods are limited to uniformly periodic repetitions.
-For non-periodic repetitions, localisation frameworks [11, 36, 49, 33] have been adapted.
+For non-periodic repetitions, temporal understanding frameworks [10, 41, 55, 37] have been adapted.
 Zhang et al.
 [70] proposed a context-aware scale-insensitive framework to count repetitions of varying scales and duration.
-Their method exhaustively searched for pairs of consecutive repetitions followed by a prediction refinement module.
+Their method exhaustively searches for pairs of consecutive repetitions followed by a prediction refinement module.
 Recent methods [14, 15, 22] have also extended image self-similarity to the temporal dimension with Temporal Self-similarity Matrices (TSM).
@@
 To count repetitions with varying speeds, Trans-RAC [22] modified TSM to use multi-scale sequence embeddings.
-For counting under poor lighting conditions, [67] used audio in tandem with video within a multi-modal framework.
+For counting under poor lighting conditions, [72] used both audio and video in a multi-modal framework.
 They selectively aggregated information from the two modalities using a reliability estimation module.
@@
 [33] also used multi-modal inputs with optical flow as an additional signal supporting RGB for detecting periodicity.
+Recent works attempt to utilise spatial [68] or temporal [73, 34] saliency for repetition counting.
 Yao et al.
-[63] proposed a lightweight pose-based transformer model that used action-specific salient poses as anchors and matched them across the video.
-However, the requirement of salient pose labels for each action limits its generalisability to unseen repetitions.
+[68] proposed a lightweight pose-based transformer model that used action-specific salient poses as anchors.
+The need for salient pose labels for each action limits generalisability to unseen repetitions.
+Zhao et al.
+[73] used a dual-branch architecture to first select repetition-relevant video segments and then attend over these frames.
+Li et al.
+[34] used a joint objective to localise and binary classify regions as (non-)repetitive.
 The above methods do not utilise the correspondences discovered by exemplar repetitions.
-Thus, do not relate variations in the action’s performance, alongside other visual variations such as viewpoint and camera motion.
-To handle such cases, we propose using action exemplars as references for VRC.
-Exemplars have previously been used in videos for action recognition tasks [17, 23, 59, 58].
-[58] used silhouette/pose exemplars for classifying action sequences into predefined action categories using a Bayes classifier.
-[59] converted training videos to visual vocabulary and used the most discriminative visual words as exemplars for detecting actions.
-These methods are limited to a predefined set of classes and will be ineffective for class-agnostic tasks such as repetition counting.
+Thus, do not relate variations in the action’s performance.
+We propose using action exemplars as references for VRC.
+Exemplars have previously been used in videos for action recognition tasks [18, 26, 64, 63][63] used silhouette/pose exemplars for classifying action sequences into predefined categories.[64] converted training videos to a visual vocabulary and used the most discriminative visual words as exemplars.
+These methods are limited to a predefined set of classes.
 To our knowledge, we are the first to use exemplars for repetition counting in videos.
 3 Every Shot Counts (ESCounts) Model
-In this section, we introduce our ESCounts model (overviewed in Figure 2).
-We formally define encoding variable length videos alongside our model’s output, in Section 3.1.
-We then introduce the attention-based decoder that corresponds the input video to exemplars in Section 3.2.
-Predictions over temporally shifted inputs are combined for a final prediction, detailed in Section 3.3.
+In this section, we introduce our ESCounts model (overviewed in Fig.
+2).
+We formally define encoding variable length videos alongside our model’s output, in Sec.
+3.1.
+We introduce the attention-based decoder that corresponds the input video to training exemplars and learnt latents in Sec.
+3.2.
+Predictions over temporally shifted inputs are then combined, detailed in Sec.
+3.3.
 Figure 2: ESCounts Model overview.
@@
 3.1 Input Encoding and Output Prediction
-We denote the full video as $\mathbf{v}$ of varying $\mathcal{T}$ length and a fixed $H\times W$ spatial resolution.
+We denote the full video as $\mathbf{v}$ of varying $\mathcal{T}$ length and fixed $H\times W$ spatial resolution.
 Segment $\mathbf{e}_{s}$ containing a single instance of the repeating action we wish to count, is selected as an exemplar.
-Exemplars are defined based on provided [[[[start, end]]]] labels of every repetition in the video [^1] For datasets where the start/end times are not available, pseudo-labels are used instead by uniformly dividing the video given the ground truth count..
+Exemplars are defined based on provided [[[[start, end]]]] labels of every repetition in the video [^1] For datasets where the start/end times are not available, pseudo-labels are used instead by uniformly dividing the video by the ground truth count..
 During training, we select one or more exemplar shots $\mathcal{S}\subseteq\{\mathbf{e}_{1},\dots,\mathbf{e}_{s}\}$.
-Each training sample is thus a combination of the query video and a set of exemplars $(\mathbf{v},\mathcal{S})$.
+Each training instance is a combination of the query video and the set of exemplars $(\mathbf{v},\mathcal{S})$.
 We tokenise and encode the video $\mathbf{v}$ from its original size $\mathcal{T}\times H\times W$ into spatiotemporal latents $\mathbf{z}_{v}$.
-Encoder $\mathcal{E}$ is applied over a fixed-size sliding window, to account for the video’s variable length.
+To account for the video’s variable length, encoder $\mathcal{E}$ is applied over a fixed-size sliding window.
 The encoded video is represented by $\mathbf{z}_{v}\in\mathbb{R}^{M\times C}$ of $M=\mathcal{T^{\prime}}H^{\prime}W^{\prime}$ spatiotemporal resolution with $C$ channels.
@@
 We add sinusoidal positional encoding to account for the relative order of these spatiotemporal latents while accommodating the variable video length.
-During training, we select exemplars $\mathcal{S}$ from either the same video or another video containing the same action label; e.g.
+For training only, we select exemplars $\mathcal{S}$ from either the same video or another video of the same action category; e.g.
 given a video containing push-up actions, we can sample exemplars from other videos showcasing the same action within the training set.
@@
 $p=0$ implies exemplars are only sampled from the same video, whereas for $p=1$ exemplars are always sampled from another video [^2] We ablate $p$ in our experiments..
-In training, exemplars are sampled randomly from the labelled repetitions of the video.
+We sample exemplars randomly from the labelled repetitions of the video.
 We use $\mathcal{E}$ to encode latent representations from each exemplar $\mathbf{e}_{s}\in\mathcal{S}$.
-We use the same encoder $\mathcal{E}$ for encoding $\mathbf{v}$ and $\mathbf{e}_{s}$ to allow correspondence.
+We use the same encoder $\mathcal{E}$ for encoding $\mathbf{v}$ and $\mathbf{e}_{s}$ to enable direct correspondence.
 We construct the ground truth density map $\mathbf{d}$ from the labelled repetitions in the video as a 1-dimensional vector.
 To match the downsampled temporal resolution of our input video $\mathcal{T^{\prime}}$, we also temporally downsample the ground-truth labels.
-The density map takes low values ($\approx 0$) at temporal locations without repetitions in the video, and high within repetitions.
+The density map takes low values ($\approx\!0$) at temporal locations without repetitions and high values within repetitions.
 We use a normal distribution $\mathcal{N}$ centred around each repetition with $(\mu_{i}=\frac{t_{s}+t_{e}}{2},\sigma)$, where $t_{e}$ and $t_{s}$ are the start and end times of each repetition $i$.
@@
 3.2 Latent Exemplar Correspondence
-Given both the encoded video $\mathbf{z}_{v}=\mathcal{E}(\mathbf{v})$ and exemplars $\mathbf{z}_{s}=\mathcal{E}(\mathbf{e}_{s})\;\forall\;\mathbf{e}_{s}\in\mathcal{S}$ latents, we use an attention-based decoder $\mathcal{D}(\mathbf{z}_{v},\mathbf{z}_{s})$ to attend to every repetition in the video $v$ that matches the encoded exemplar.
+Given both the encoded video $\mathbf{z}_{v}=\mathcal{E}(\mathbf{v})$ and exemplars $\mathbf{z}_{s}=\mathcal{E}(\mathbf{e}_{s})\;\forall\;\mathbf{e}_{s}\in\mathcal{S}$, we use an attention-based decoder $\mathcal{D}(\mathbf{z}_{v},\mathbf{z}_{s})$ to learn a correspondence between every repetition in the video $v$ and the encoded exemplar.
 Decoder $\mathcal{D}$ takes the encoded video $\mathbf{z}_{v}$ as input and predicts the location of every repetition in the video.
@@
 Cross-attention Blocks.
-We explore the similarity between representations of exemplars and the query video to predict the corresponding locations of repetitions that match the exemplar.
-Thus, inspired by [10], we use cross-attention to relate exemplar and video latents.
+We explore the similarity between exemplars and query video representations to predict the corresponding locations of repetitions that match the exemplar.
+Thus, inspired by [38], we use cross-attention to relate exemplar and video encodings.
 We define $L$ cross-attention blocks.
@@
 We note that for the first layer, $\mathbf{z}_{1}=\mathbf{z}_{v}$.
-Then we relate exemplar and video information by cross-attending $\text{CA}(\cdot)$ video and exemplar latents.
+We then relate exemplar and video by Cross-Attending $\text{CA}(\cdot)$ video and exemplar encodings.
 The block’s initial self-attention operation is formulated as:
@@
 where $\text{LN}(\cdot)$ is Layer Normalisation.
+It is essential to self-attend across the video first to capture the features of the repeated actions within the video, and enforce feature correspondence between repetitions.
 […]
@@
 where $\mathcal{S}$ is the set of exemplars selected and $\mathbf{z}_{s}$ is the latent for the $s^{\text{th}}$ exemplar.
-We also want to learn representations for repeating motions to estimate repetitions without explicitly providing exemplars.
+We also want to learn repeating motions to estimate repetitions without explicitly providing exemplars.
 We thus define a learnable latent $\mathbf{z}_{0}$ to cross-attend $\mathbf{z}_{v}$.
-At each training iteration, we select exemplars from $\{\mathbf{z}_{0},\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$ and perform CA with $\mathbf{z}_{0}$ or $\{\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$.
-At inference, we use $\mathbf{z}_{0}$.
-We use a multi-layer perceptron MLP on the exemplar-fused latents $\mathbf{z}^{\prime\prime}_{l}$.
+At each training step, we select exemplars from $\{\mathbf{z}_{0},\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$ and perform CA with $\mathbf{z}_{0}$ or $\{\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$.
+Importantly, at inference, we use only $\mathbf{z}_{0}$.
+We obtain the cross-attention blocks’ output, defined as $\mathbf{z}_{l+1}\in\mathbb{R}^{M\times C}$, with a Multi-Layer Perceptron MLP on the exemplar-fused latents $\mathbf{z}^{\prime\prime}_{l}$.
 $\mathbf{z}_{l+1}=\text{MLP}(\text{LN}(\mathbf{z}^{\prime\prime}_{l}))+\mathbf{z}^{\prime\prime}_{l}\,\;\forall\,l\in\,\{1,\dots,L\}$ (4)
-The cross-attention blocks’ output is defined as $\mathbf{z}_{L}\in\mathbb{R}^{M\times C}$.
 Window Self-attention Blocks.
 We explore the spatio-temporal inductive bias within the self-attention blocks.
-Each latent thus attends locally, to its spatio-temporal neighbouring tokens, over $L^{\prime}$ window self-attention $\text{WSA}(\cdot)$ [35] layers.
+For this, each latent attends locally to its spatio-temporal neighbouring tokens, over $L^{\prime}$ Window Self-Attention $\text{WSA}(\cdot)$ [40] layers.
 We denote $\forall\,l\in\,\{L+1,\dots,L+L^{\prime}\}$ :
@@
 where WSA is window self-attention. Note that following [40] windows are shifted at each layer to account for connections across different windows.
-The output of the window self-attention blocks is of size $\mathbf{z}_{L+L^{\prime}}\in\mathbb{R}^{M\times C}$ In turn, $\mathbf{z}_{L+L^{\prime}}$ encodes repetition-relevant features over space and time.
-Thus, we use $\mathbf{z}_{L+L^{\prime}}$ to predict density map $\tilde{\mathbf{d}}$ for the occurrences of the target repeating action over time.
+The output of the WSA blocks is of size $\mathbf{z}_{L+L^{\prime}}\in\mathbb{R}^{M\times C}$ In turn, $\mathbf{z}_{L+L^{\prime}}$ encodes repetition-relevant features over space and time and is used to predict density map $\tilde{\mathbf{d}}$ for the occurrences of the target repeating action over time.
 We use a fully connected layer to project the latent to a 1-channel vector, i.e.
 MLP: ${\mathbb{R}^{M\times C}\rightarrow\mathbb{R}^{M}}$.
-We then pool the spatial resolution $H^{\prime}W^{\prime}$ resulting to predicted density map $\tilde{\mathbf{d}}\in\mathbb{R}^{\mathcal{T}^{\prime}}$.
+We then vectorise the spatial resolution $H^{\prime}W^{\prime}$ whilst maintaining $\mathcal{T}^{\prime}$ resulting to the predicted density map $\tilde{\mathbf{d}}\in\mathbb{R}^{\mathcal{T}^{\prime}}$.
 Training Objective.
@@
 We train with augmentations of the start time whilst at inference, we use an ensemble of time-shift augmentations for a more robust estimation.
-We use multiple overlapping sequences as shown in Figure 4 and combine the predicted density maps over $\mathbf{K}$ shifted start/end positions.
-We obtain the final predicted density map by temporally aligning and averaging the predictions.
-$\tilde{\mathbf{d}}_{t}=\frac{1}{|\mathbf{K}|}\underset{k\in\mathbf{K}}{\sum}\tilde{\mathbf{d}}^{k}_{t+\epsilon_{k}}$ (7)
-where $\epsilon_{k}$ is the shifting for each $k\in\mathbf{K}$.
+We use multiple overlapping sequences as shown in Fig.
+4 and combine the predicted density maps over $\mathbf{K}$ shifted start/end positions.
+We obtain the final predicted density map by temporally aligning and averaging the predictions; $\tilde{\mathbf{d}}_{t}=\frac{1}{|\mathbf{K}|}\underset{k\in\mathbf{K}}{\sum}\tilde{\mathbf{d}}^{k}_{t+\epsilon_{k}}$, where $\epsilon_{k}$ is the shifting for each $k\in\mathbf{K}$.
 4 Experiments
@@
 4.3.
-We demonstrate ESCounts’ ability to utilise exemplars at inference, in Sec.
-4.4.
+For all results, we only report zero-shot counting during inference.
+In Sec.
+4.4, we evaluate ESCounts’ when exemplars are available during inference.
 4.1 Experimental Setup
 Datasets. We evaluate our method on a diverse set of VRC datasets.
-RepCount [20] contains videos of workout activities of varying repetition durations and appearances.
-Annotations include count labels alongside start and end timestamps per repetition.
-We use the publicly available Part-A of the dataset with $758$, $131$, and $152$ videos for train, val, and test respectively.
+RepCount [22] contains videos of workout activities with varying repetition durations.
+Annotations include counts alongside start and end times per repetition.
+We use the publicly available Part-A with $758$, $131$, and $152$ videos for train, val, and test respectively.
+Additionally, we use the provided open set split with 70% categories for train, 10% for val, and 20% for testing.
 We tune the hyperparameters on the val set and report our results on the test set.
-Countix [16] is a subset of Kinetics [9] containing videos of repetitive actions and their associated counts.
-It is the largest VRC dataset with $4588$, $1450$, and $2719$ videos for train, val, and test respectively.
-Count annotations are provided per video, without individual repetition start-end times.
-We use pseudo-repetition annotations by dividing the videos into uniform segments based on the ground truth count, as Countix does not have many pauses or interruptions between counts.
-We use the pseudo-labels for estimating the density maps, without additional annotations, to compare directly to other methods.
-UCFRep [65] is a subset of UCF-101 [51] consisting of 420 videos for train and 106 videos for val from 23 categories.
+Countix [15] is a subset of Kinetics [9] containing videos of repetitive actions with $4,588$, $1,450$, and $2,719$ videos for train, val, and test respectively.
+Counts are provided without individual repetition start-end times.
+Countix does not have many pauses or interruptions between counts.
+Thus, we define pseudo-repetition annotations by dividing videos into uniform segments based on the ground truth count.
+The pseudo-labels are used to estimate the density maps without additional annotations, to compare directly to other methods.
+UCFRep [70] is a subset of UCF-101 [57] consisting of 420 train and 106 val videos from 23 categories with counts and annotations of start and end times.
 Following [70, 33], we report our results on the val split as no test set is available.
-Like RepCount, it includes annotations of start and end timestamps for each repetition.
 Implementation Details.
-Unless specified otherwise, we use VideoMAE-VIT-B [55] as our encoder $\mathcal{E}$ with weights pre-trained on Kinetics-400 [9].
+Unless specified otherwise, we use MAE-pretrained ViT-B [17] as our encoder $\mathcal{E}$ with Kinetics-400 [9] weights.
 We sample frames from variable-length videos every 4 frames using a sliding window of $64$ frames.
@@
 Exemplars are sampled uniformly with $16$ frames between the start and end of a repetition.
-The encoder is frozen with only the decoder and the zero-shot latent $\mathbf{z}_{0}$ being trained.
-We use $L=2$ and $L^{\prime}=3$.
-We ablate $L$ and $L^{\prime}$ in the supplementary material.
-We train our model on a single Tesla V100 for 300 epochs with a batch size of 1, to deal with variable-length videos, and accumulate gradients over 8 batches.
-We use an initial learning rate of $55e-$ 5 and decay it every $60$ epochs by $0.8$, and a weight decay of $25e-$ 2.
-For each training instance, we randomly set the number of exemplars $|\mathcal{S}|\sim\{0,1,2\}$ and sample $\mathcal{S}$ exemplars.
+The encoder is frozen and we only train the decoder and zero-shot latent $\mathbf{z}_{0}$.
+We use $L=2$ and $L^{\prime}=3$ ablating this choice in Sec.
+7 of the appendix.
+We train for 300 epochs on a single Tesla V100 with a batch size of 1, to deal with variable-length videos, accumulating gradients over 8 batches.
+We use $5e^{-2}$ weight decay and a learning rate of $5e^{-5}$ with decay by $0.8$ every $60$ epochs.
+Per training instance, we randomly set the number of exemplars $|\mathcal{S}|\!\sim\!\{0,1,2\}$ and sample $\mathcal{S}$ exemplars.
 We set the chance of sampling exemplars from a different video to $p=0.4$.
-At inference, we only use the learnt latent to predict repetition counts. We aggregate predictions over $|\mathbf{K}|=4$ sequences.
+Only the learnt latent are used at inference to predict repetition counts. We aggregate predictions over $|\mathbf{K}|=4$ sequences.
 Evaluation Metrics.
-Following previous repetition counting works [20, 16, 67], we use Mean Absolute Error (MAE) and Off-By-One accuracy (OBO) as evaluation metrics, calculated as Eqs.
-8 and 9 respectively.
+Following previous VRC works [22, 15, 72], we use Mean Absolute Error (MAE) and Off-By-One accuracy (OBO) as evaluation metrics, calculated as Eqs.
+7 and 8 respectively.
 Inspired by image counting methods [38, 3], we introduce Root-Mean-Square-Error (RMSE) in Eq.
 9 for VRC providing a more robust metric for diverse counts compared to MAE’s bias towards small counts.
-With continuous improvements in repetition counting methods, we also report the off-by-zero accuracy (OBZ) in Eq.
+We also report the off-by-zero accuracy (OBZ) in Eq.
 10 as a tighter metric than the corresponding OBO for precise counts.
@@
 In Tab.
-2, we compare our method, ESCounts, to prior methods on the three datasets.
-We also provide results on the same backbone as the best-performing method on each dataset, for fair and direct comparison to previous works.
+2, we compare ESCounts, to prior methods on the three datasets.
+We provide results on the same backbone as the best-performing method on each dataset, for fair and direct comparison to previous works.
 Table 1: Comparison of VRC methods.
@@
 denotes results reproduced using provided checkpoints.
-denotes inhouse retraining using published codes.
+denotes inhouse re-training using published codes.
 Grayed rows in (c) represent methods that finetune the encoder.
@@
 (a) RepCount
+benchmark open set
 […]
-RepNet [16] R2D50 - 0.995 - 0.013
-TransRAC [20] VSwinT  9.130 0.443  0.085 0.291
-MFL [30$\dagger$ VSwinT - 0.384 - 0.386
-ESCounts VSwinT 6.905 0.298  0.1830.183\;0.183\;0.183 0.403
-ESCounts VMAE  $\;\mathbf{4.455}\;$ $\mathbf{0.213}$  $\;\mathbf{0.245}\;$ $\mathbf{0.563}$
+RepNet [15] R2D50 - 0.995 - 0.013 - -
+TransRAC [22] SwinT  9.130 0.443  0.085 0.291 0.625 0.204
+MFL [33$\dagger$ SwinT - 0.384 - 0.386 - -
+DeTRC [34] ViT-B - 0.262 - 0.543 - -
+SkimFocus [73] SwinB - 0.249 - 0.517 - -
+ESCounts SwinT 6.905 0.298 $0.183$ 0.403 - -
+ESCounts ViT-B $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$ $\mathbf{0.436}$ $\mathbf{0.519}$
 (b) Countix
@@
 RepNet [15] R2D50 - 0.364 - 0.697
-Sight & Sound [67$\dagger$ R(2+1)D18 - 0.307 - 0.511
+Sight & Sound [72$\dagger$ R(2+1)D18 - 0.307 - 0.511
 ESCounts R(2+1)D18 3.536 0.293 0.286 $\mathbf{0.701}$
-ESCounts VMAE $\mathbf{3.029}$ $\mathbf{0.276}$ $\mathbf{0.319}$ 0.673
+ESCounts ViT-B $\mathbf{3.029}$ $\mathbf{0.276}$ $\mathbf{0.319}$ 0.673
 (c) UCFRep
@@
 Context (F) [70] RX3D101 5.761 0.653 0.143 0.372
-TransRAC [20] VSwinT - 0.640 - 0.324
-MFL[30$\dagger$ RX3D101 - 0.388 - 0.510
+TransRAC [22] SwinT - 0.640 - 0.324
+MFL[33$\dagger$ RX3D101 - 0.388 - 0.510
 ESCounts RX3D101 2.004 0.247 0.343 0.731
-ESCounts VMAE $\mathbf{1.972}$ 0.216 0.381 0.704
+ESCounts ViT-B 1.972 0.216 0.381 0.704
 Context [70] RX3D101 2.165 0.147 $\mathbf{0.452}$ 0.790
-Sight & Sound [67$\dagger$ R(2+1)D18 - $\mathbf{0.143}$ $\mathbf{0.800}$
-Table 2: Cross-dataset generalisation scores. Arrows $X\rightarrow Y$ denote train dataset $X$ and test dataset $Y$. Results obtained using provided checkpoints are denoted with .
-RepCount $\rightarrow$ UCFRep RepCount $\rightarrow$ Countix
+Sight & Sound [72$\dagger$ R(2+1)D18 - $\mathbf{0.143}$ $\mathbf{0.800}$
+Table 2: Cross-dataset generalisation. Arrows denote train $\rightarrow$ test datasets. Results with provided checkpoints are denoted with .
+RepCount $\rightarrow$ Countix RepCount $\rightarrow$ UCFRep
 […]
-RN [16] - 0.998 - 0.009 - - - -
-TRAC [20] 6.701 0.640 0.087 0.324 6.867 0.593 0.132 0.364
-MFL [30] - 0.523 - 0.350 - - - -
-ESCounts $\mathbf{3.536}$ $\mathbf{0.317}$ $\mathbf{0.219}$ $\mathbf{0.571}$ $\mathbf{4.429}$ $\mathbf{0.374}$ $\mathbf{0.185}$ $\mathbf{0.521}$
-Tab.
-2 - RepCount.
-We demonstrate that our ESCounts model outperforms recent methods [20, 30] with the same Swin Tiny (VSwinT) [35] encoder, by $-0.15$ in MAE and $+0.11$ in OBO.
-Our approach also improves upon [30] which uses optical flow and video in tandem, by margins of $-0.09$ in MAE and $+0.02$ in OBO.
-The largest improvements are observed with VideoMAE (VMAE) surpassing [20] on the introduced OBZ by $+0.16$ and reducing RMSE by $-4.68$.
-Tab.
-2 - Countix.
-Compared to the state-of-the-art [16, 67] our ESCounts consistently outperforms other models with the same R(2+1)D18 encoder.
+RepNet [15] - - - - - 0.998 - 0.009
+TransRAC [22] 6.867 0.593 0.132 0.364 6.701 0.640 0.087 0.324
+MFL [33] - - - - - 0.523 - 0.350
+SkimFocus [73] - - - - - 0.502 - 0.391
+DeTRC [34] - - - - - 0.543 - 0.418
+ESCounts $\mathbf{4.429}$ $\mathbf{0.374}$ $\mathbf{0.185}$ $\mathbf{0.521}$ $\mathbf{3.536}$ $\mathbf{0.317}$ $\mathbf{0.219}$ $\mathbf{0.571}$
+RepCountTab.
+2 shows that ESCounts outperforms recent methods [22, 33, 34, 73].
+Compared to the baseline [22], we improve OBZ by $+0.16$ and reduce RMSE by $-4.68$.
+We test on two backbones - SwinT [40] used in [22, 33] and ViT-B used in [34].
+On the same SwinT backbone, our approach outperforms [33], which uses optical flow and video in tandem, by margins of $-0.09$ MAE and $+0.02$ OBO, showcasing ESCounts’ ability to learn repeating motions implicitly.
+With a ViT-B backbone, we outperform [34] by $-0.05$ MAE and $+0.02$ OBO.
+We additionally compare ESCounts on the open set setting of RepCount-A, with non-overlapping action categories between train and test sets.
+ESCounts outperforms [22] significantly with $-0.19$ MAE and $+0.32$ OBO.
+Note that recent works do not report on this more challenging setup.
+Countix.
+Compared to the state-of-the-art [15, 72] in Tab.
+2, our ESCounts consistently outperforms other models with the same R(2+1)D18 encoder.
 Our video-only model surpasses the audio-visual model in [72] by $+0.19$ OBO.
-Further improvements on the RMSE, MAE, and OBZ are observed with VMAE.
-Tab.
-2 - UCFRep.
-Across approaches with frozen encoders ESCounts improve the top-performing [30] by $+0.19$ OBO and $-0.17$ MAE.
-Our method does not outperform [67, 65] that fine-tune their encoders on the training set of UCFRep which is advantageous given the dataset’s size as also noted in [30].
-We showcase this with Context (F) trained from the publicly available code of [65] with a frozen encoder and result in a significant drop in performance with +0.51 MAE and -0.42 OBO.
-Despite this, we achieve strong results.
-Figure 5: Scatter plot, instances, and density maps from RepCount, Countix and UCFRep.
+Further improvements on the RMSE, MAE, and OBZ are observed with ViT-B.
+UCFRep.
+Compared to methods with frozen encoders in Tab.
+2, ESCounts with ViT-B improves the previous SoTA by $+0.19$ OBO and $-0.17$ MAE and outperforms [33] on the same RX3D101 backbone by $+0.22$ OBO.
+Our method does not outperform [72, 70] that fine-tune their encoders on UCFRep.
+As noted in [33] this is advantageous given the dataset’s size.
+We show this experimentally by reporting Context (F) trained from the available code of [70] with a frozen encoder, resulting in a significant performance drop with +0.51 MAE and -0.42 OBO.
+In all directly comparable results, ESCounts achieves stronger results.
+Figure 5: RepCount, Countix, and UCFRep scatter plot, instances, and density maps.
 The dotted diagonal denotes correct predictions.
-We compare ESCounts against TransRAC [20] on Repcount and Context [65] on UCFRep.
-Examples a--m are marked with corresponding action class and count predictions.
-We add the Ground Truth (GT) and Predicted (P) density maps for each instance.
-For Countix, pseudo-labels are shown as GT, as no start-end times are given.
+We compare ESCounts against TransRAC on Repcount and Context on UCFRep.
+Action classes and count predictions are shown for each instance.
+We add the Ground Truth (GT) and Predicted (P) density maps per instance.
+Pseudo-labels are shown as GT for Countix.
 Qualitative Results.
 In Fig.
-5 we visualise predicted to ground truth counts.
+5 we visualise predicted to ground truth counts as scatter plots.
 For RepCount and UCFRep, we select [22] and [70] as respective baselines and use their publicly available checkpoints444[33] was not used as a baseline as the code is not public.
 The publicly available checkpoints on Countix obtained lower results than originally reported..
-ESCounts accurately predicts the number of repetitions for a wide range of counts, with most predictions being close to the diagonal.
+ESCounts accurately predicts the number of repetitions for a wide range of counts, with most predictions being close to the ground truth i.e.
+the diagonal.
 Though predictions from both the baseline and ESCounts are close to the ground truth in low counts, they significantly diverge in high counts.
-We additionally visualise specific examples with sample frames from these videos.
-ESCounts is robust to the magnitude of counts, with accurate predictions over low count (a,b,g,k,l) and high count (d,i,m) examples.
+We visualise specific examples and their density maps.
+ESCounts is robust to the magnitude of counts, with accurate predictions over low (a,b,g,k,l) and high (d,i,m) count examples.
 In cases of over- and under-predictions; e.g.
@@
 As shown by the density maps, ESCounts can also localise the repetitions.
-For Countix, because the model was trained on pseudo labels, even though ESCounts can predict accurate counts, it struggles to localise the repetitions at times.
-We further investigate the localisation capabilities in the supplementary.
+For Countix, even though ESCounts can predict accurate counts, as the model was trained on pseudo labels, it struggles to localise some of the repetitions.
+We investigate the localisation capabilities in Sec.
+8 of the appendix.
 Cross-dataset Generalisation.
-Following prior works [20, 30], we also test the generalisation capabilities of our method.
-We use our model trained on RepCount and evaluate it on the test sets of UCFRep and Countix.
-As shown in Tab.
-2, our method achieves the best results surpassing [30] by $-0.21$ in MAE and $+0.22$ in OBO in UCFRep.
-We also explore cross-dataset generalisation for Countix, and select [20] as a baseline with the public checkpoint.
-We outperform [20] by significant margins across metrics, showcasing the strong ability of ESCounts to generalise across datasets.
+Following [22, 33], we test the generalisation capabilities of our method in Tab.
+2.
+We use ESCounts trained on RepCount and evaluate on the Countix and UCFRep test sets.
+For Countix, we outperform the baseline [22] by significant margins across metrics.
+For UCFRep, our method surpasses [33] by $-0.21$ in MAE and $+0.22$ in OBO.
+ESCounts in this setting still outperforms [15, 22, 33, 34] trained on UCFRep in Tab.
+2, showcasing the strong ability of ESCounts to generalise to unseen actions.
 4.3 Ablation Studies
-In this section, we conduct ablation studies on RepCount [20] using VMAE as the encoder.
-We study the impact of exemplars by replacing cross- with self-attention and vary the number of exemplars used during training.
-We evaluate the sensitivity of our method to the exemplar sampling probability $p$.
-We ablate $\sigma$ in density map estimation and analyse the impact of time-shift augmentations and our loss function’s components.
+In this section, we conduct ablation studies on RepCount [22] using ViT-B as the encoder.
+We study the impact of exemplars by replacing cross- with self-attention and varying the number of training exemplars.
+We evaluate the sensitivity of our method to the exemplar sampling probability $p$, density map $\sigma$, the impact of time-shift augmentations, and the components of our objective.
 Table 3: Ablations on RepCount over different ESCounts settings.
@@
 $|\mathcal{S}|\sim\{0,2\}$ 4.601 0.226 0.239 0.550
-$|\mathcal{S}|\sim\{0,1,2\}$ $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
+$|\mathcal{S}|\sim\{0,1,2\}$ $\mathbf{4.455}$ $\mathbf{0.213}$ 0.245 $\mathbf{0.563}$
+$|\mathcal{S}|\sim\{0,1,2,3\}$ 4.497 0.215 $\mathbf{0.246}$ 0.560
+$|\mathcal{S}|\sim\{0,1,2,3,4\}$ 4.482 0.215 0.240 0.559
 (c) Exemplar sampling
@@
 3, using self-attention (SA-only) performs significantly worse than our proposed ESCounts.
-Cross-attending exemplars decrease the RMSE/MAE by $-1.20$ and $-0.06$ whilst improving OBZ and OBO by $+0.10$ and $+0.09$, thus emphasising the benefits of exemplar-based VRC.
+Cross-attending exemplars decrease the RMSE/MAE by $-1.20$ and $-0.06$ whilst improving OBZ and OBO by $+0.10$ and $+0.09$, respectively.
+This emphasises the benefits of exemplar-based VRC.
 How many exemplars to sample?
-Since we use a varying number of exemplars whilst training, we also experiment by varying $|\mathcal{S}|$ from $0-2$ in Tab.
+A varying number of training exemplars $|\mathcal{S}|$ is used in Tab.
 3.
-In $|\mathcal{S}|=0$ we train only with the learnt zero-shot latent.
-Training with $|\mathcal{S}|\sim\{0,1,2\}$ provides the best zero-shot scores in inference.
-We believe training with a larger and more diverse set of exemplars helps the model to learn a general zero-shot token that can capture general information about repetitions.
-Due to memory constraints, we could not train with more than 2 exemplars.
+For $|\mathcal{S}|=0$, we train only the zero-shot latent $\mathbf{z}_{0}$ alongside the model’s parameters.
+Training with $|\mathcal{S}|\sim\{0,1,2\}$ provides the best zero-shot scores at inference with our method efficiently learning to generalise by attending to only a few exemplars.
+The inclusion of more exemplars saturates performance.
 How to sample exemplars?
-We analyse the impact of sampling exemplars from other videos, from the same or a different action class.
-As expected, Tab.
-3 shows that keeping the same action category for both exemplars and query video performs the best, as ensuring the same action semantics between exemplars and query video helps to learn their correspondence.
+In Tab.
+3 we analyse the impact of sampling exemplars from the same or other training videos.
+As expected, keeping the same action category for both exemplar and query videos performs the best, as ensuring the same action semantics between exemplars and query video helps to learn correspondence.
 In this table, we used sampling probability $p=0.4$.
-We vary the sampling probability from other videos of the same underlying action $p$ in Tab.
-3.
-For $p=0.0$ exemplars are sampled exclusively from the query video, whilst for $p=1.0$ exemplars are sampled solely from other videos of the same class.
-The best performance was observed with $p=0.4$, showcasing sampling from the same video to be critical for learning.
+In Tab.
+3, we vary the sampling probability from other videos of the same underlying action $p$.
+For $p=0.0$, exemplars are sampled exclusively from the query video, whilst for $p=1.0$, exemplars are sampled solely from other videos of the same class.
+The best performance was observed with $p=0.4$, showcasing that the visual characteristics of exemplars from the same video are critical for VRC compared to class semantics.
 […]
@@
 What’s the impact of time-shift augmentations?
-As predictions are aggregated over $|\mathbf{K}|$ density maps by time-shifting the video input.
+Predictions are aggregated over $|\mathbf{K}|$ density maps by time-shifting the video input.
 As shown in Tab.
@@
 The combined objective helps performance for diverse counts across all metrics.
-How close are predictions to the ground truth?.
-We further relax the off-by metrics to Off-By-N in Figure 6(a) to visualise the proximity of predictions to the ground truth.
+How close are predictions to the ground truth?
+We further relax the off-by metrics to Off-By-N in Fig.
+6(a) to visualise the proximity of predictions to the ground truth.
 Overall, 84% of predictions are within ±$\pm 3$ of the actual count.
-What is the performance per action category?.
-In Figure 6(b), we plot the OBZ per action class.
-ESCounts performs fairly uniformly across all classes with the best performing categories being pommelhourse and squat.
+What is the performance per action category?
+In Fig.
+6(b), we plot the OBZ per action class.
+ESCounts performs fairly uniformly across all classes with the best-performing categories being pommelhourse and squat.
 How does performance differ across counts, repetition lengths, and video durations?
@@
 Predicting density maps is more challenging for short repetitions.
-However, ESCounts can still correctly predict counts across repetition lengths as shown by Figures 6(i) and 6(j).
+However, ESCounts can still correctly predict counts across repetition lengths as shown by Figs.
+6(i) and 6(j).
 We also group videos by duration into XS=(8.0-11.0)s, S=(11.0, 26.0)s, M=(26.0, 33.9)s, L=(33.9-45.9)s and XL=(45.9-68.0)s.
-From Figures 6(k), 6(l), 6(m) and 6(n), counting repetitions from longer videos is more challenging.
+From Figs.
+6(k), 6(l), 6(m) and 6(n), counting repetitions from longer videos is more challenging.
 Table 4: Number of shots at inference. We test using exemplars from the same video or a different video of the same action class from the train set.
@@
 $\mathbf{4.351}$ $\mathbf{0.206}$ 0.250 $\mathbf{0.596}$ $\mathbf{1.855}$ $\mathbf{0.198}$ $\mathbf{0.420}$ 0.723
-4.4 Inference with More Shots
-up to this point, only the learnt zero-shot latent was used for inference.
-Prior works in object counting [10, 38] have reported results using exemplars (i.e.
-object crops) in inference.
-While this is not comparable to other works in VRC, we can assess our method’s ability to utilise exemplars during inference in Tab.
+4.4 Multi-Shot Inference
+We use learnt latents for exemplar-free inference.
+Prior object counting [38, 43] report results with exemplars (i.e.
+object crops) at inference.
+While this is not comparable to other VRC works, we can assess our method’s ability to utilise exemplars during inference in Tab.
 4.
 Video exemplars steadily improve performance as the number of exemplars increases.
-As our model cross-attends exemplars in parallel, we can train with $0-2$ exemplars, and use $>2$ exemplars during inference.
-We show comparable results when sampling exemplars from the test video vs exemplars from the training set videos showcasing the same action category.
-Combined with a classifier, one can envisage an approach that classifies the action then sources exemplars from the training set to assist counting during inference.
+Our model cross-attends exemplars in parallel, training with $0\!\!-\!\!2$ exemplars, and can even use $>\!2$ exemplars at inference.
+We show comparable results when sampling exemplars from the test video or training videos with the same action category.
+Combined with a classifier, a closed-set approach can be envisaged that classifies the action and then sources exemplars from the training set to assist counting during inference.
 5 Conclusion
@@
 Extensive evaluation on RepCount, Countix, and UCFRep demonstrates the merits of ESCounts achieving state-of-the-art results on the traditional MAE and OBO metrics and the newly introduced RMSE and OBZ.
-Further analysis showed that the exemplars do not need to be from the same video as long as they belong to the same action class.
+We provide detailed analysis and ablations of our method, highlighting the importance of training with exemplars and time-shift augmentations.
 The diversity of these exemplars is an aspect for future exploration.
-Acknowledgements
-This work uses publicly available datasets and annotations to produce all results and ablations.
+Acknowledgements.
+This work uses publicly available datasets and annotations for results and ablations.
 Research is supported by EPSRC UMPIRE (EP/T004991/1).
 S.
-Sinha is supported by EPSRC DTP (Doctoral Training Program) studentship.
+Sinha is supported by EPSRC DTP studentship.
 Appendix
-Code is made available anonymously from: https://github.com/sinhasaptarshi/EScounts. We provide the full code, checkpoint as well as a demo with a couple of videos to try our inference.
-In this supplementary, we explain the attached qualitative video in Sec.
+Code is made publicly available at: https://github.com/sinhasaptarshi/EveryShotCounts. The repository contains the full train and evaluation code and a demo for inference with a few videos.
+In the following sections, we provide more qualitative results in Sec.
 6.
@@
 8.
-Finally, we analyse our improvement over baseline by studying groups of videos based on their length, number of repetitions, and counts in Sec.
-9.
-6 Qualitative Video and Extended Figure
-We attach a compilation video (1380.mp4) showcasing our method’s abilities in Video Repetition Counting (VRC).
-For a diverse set of 11 videos from all 3 datasets, we show the video alongside synchronised ground truth and ESCounts predicted density maps.
-For each example, we note which test set it is from.
-Additionally, we extend Fig. 5 in the main paper with additional examples in Fig. 8.
-7 Further Ablations
-We extend the ablation studies in  Sec. 4.3 by reporting results over different $L$ and $L^{\prime}$. Additionally, we analyse the impact of windowed-self attention on the performance of ESCounts.
-(o) RepCount
-(p) Countix
-Figure 7: Additional qualitative results.
-(a) UCFRep
-Figure 8: Additional qualitative results (continued).
+We then compare VRC to Temporal Action Segmentation (TAS) in Sec.
+9 demonstrating distinctions between the two tasks.
+Additionally, following the release of the recent egocentric video counting dataset OVR-Ego4D [16], we train and evaluate ESCounts on this newly introduced dataset demonstrating the effectiveness of our method for egocentric counting in Sec.
+10.
 Table 5: Impact of $L$.
@@
 full 5.011 0.227 0.221 0.533
+6 Qualitative Video and Extended Figure
+We provide a compilation of videos on our website https://sinhasaptarshi.github.io/escounts/ showcasing our method’s Video Repetition Counting (VRC) abilities over a diverse set of 20 videos from all 3 datasets.
+Videos are shown alongside synchronised ground truth and predicted density maps.
+The test set from which each video is sampled is also shown.
+We additionally extend Fig. 5 in the main paper with more examples from all datasets in Fig. 8.
+7 Further Ablations
+We extend the ablations in Sec. 4.3, report results over different $L$ and $L^{\prime}$, and analyse the impact of windowed-self attention on the performance of ESCounts.
+(o) RepCount
+(p) Countix
+Figure 7: Additional qualitative results.
+(a) UCFRep
+Figure 8: Additional qualitative results (continued).
 Impact of $L$.
-We ablate the value of $L$ i.e.
+We ablate $L$ i.e.
 the number of layers in the cross-attention block.
@@
 $L^{\prime}$ is the number of windowed self-attention layers in the self-attention block.
-$L^{\prime}=3$ gives the best results on all the metrics.
+$L^{\prime}=3$ gives the best results across all the metrics.
 Similarly, increasing or decreasing $L^{\prime}$ drops performance gradually.
 Self-attention vs Windowed Self-attention.
-Motivated by [35], we use windowed self-attention for the self-attention block in ESCounts.
-Given spatio-temporal tokens $\mathcal{T}^{\prime}\times H^{\prime}\times W^{\prime}\times C$, windowed self-attention computes multi-headed attention for each token within the immediate neighbourhood using 3D shifted windows of size $t^{\prime}\times h^{\prime}\times w^{\prime}$, where $t^{\prime}<\mathcal{T}^{\prime}$, $h^{\prime}<H^{\prime}$ and $w^{\prime}<W^{\prime}$.
-Here we ablate for various values $(t^{\prime},h^{\prime},w^{\prime})$ in Tab.
+Motivated by [40], we use windowed self-attention for the decoder self-attention blocks.
+Given spatio-temporal tokens $\mathcal{T}^{\prime}\times H^{\prime}\times W^{\prime}\times C$, windowed self-attention computes multi-headed attention for each token within the immediate neighbourhood using 3D shifted windows of size $t^{\prime}\times h^{\prime}\times w^{\prime}$, where $t^{\prime}\leq\mathcal{T}^{\prime}$, $h^{\prime}\leq H^{\prime}$ and $w^{\prime}\leq W^{\prime}$.
+We ablate on various $(t^{\prime},h^{\prime},w^{\prime})$ values in Tab.
 7.
-Note that $t^{\prime}=\mathcal{T}^{\prime}$, $h^{\prime}=H^{\prime}$ and $w^{\prime}=W^{\prime}$ is using standard self-attention over all the embeddings, where each token is attended with every token and is denoted by full.
+Note that for $t^{\prime}=\mathcal{T}^{\prime}$, $h^{\prime}=H^{\prime}$, and $w^{\prime}=W^{\prime}$ denoted as full, standard self-attention is used where each token attends to every token.
 As shown, the best performance is obtained with window size $(4,7,7)$, demonstrating the importance of attending to tokens in immediate spatio-temporal neighbourhoods only.
-We found that the results are most sensitive to the value of $t^{\prime}$.
-Performance steadily drops as the value of $t^{\prime}$ changes.
+We found variations in the value of $t^{\prime}$ to have the largest performance impact with decreases as the value of $t^{\prime}$ changes.
 Sampling Rate for Encoding.
-As stated in the implementation details, we construct our input by sampling every four frames from the video.
+As stated in the implementation details, we sample every four frames from the video to form the encoder inputs.
 We ablate the impact of the sampling rate in Tab.
 8.
-From the results, we observe that denser sampling is key for robust video repetition counting.
-Reducing the sampling rate steadily decreases performance.
+As shown, denser sampling is key for robust video repetition counting.
+Reducing the sampling rate steadily decreases performance as relevant parts of repetitions may be missed.
 Table 8: Impact of sampling rate
@@
 32 6.562 0.346 0.156 0.444
-Table 9: OBO, parameters, and training and inference speeds on UCFRep. Mtreics obtained by the public available codebase of [65] are denoted with .
+Table 9: OBO, parameters, and training and inference speeds on UCFRep. Metrics obtained by the public available codebase of [70] are denoted with .
 […]
@@
 However, this performance is achieved by having more trainable parameters, as [70, 72] finetune the encoders on the target dataset.
-We use the provided codebase from [65] and benchmark the average number of iterations per second for a full forward and backward pass over the entire training set alongside inference-only average times on the test set.
+We use the provided codebase from [70] and benchmark the average number of iterations per second for a full forward and backward pass over the entire training set.
+Additionally, we report inference-only average times on the test set.
 We use the same experiment set-up described in Sec.
-4.1 and report speeds in Table 9.
-Training ESCounts is  8 times faster.
-Interestingly, even during inference ESCounts is  12 times faster than Context [65] which uses iterative processing.
+4.1 and report speeds in Tab.
+9.
+Training ESCounts is $\sim$8$\times$ faster.
+Interestingly, ESCounts maintains its efficiency even during inference with $\sim$12$\times$ faster times than Context [70] which uses iterative processing.
 Note that [72] could not be used for this analysis as their code for training with UCFRep is not publicly available.
@@
 ESCounts 38.83 38.64 38.07 37.44 35.82 33.43 30.76 27.52 20.85 33.48
-8 Repetition Localisation
-Figure 9: Explaining localisation metric $\mathcal{J}$.
+Figure 9: Localisation metric $\mathcal{J}$.
 We identify local maxima in $\tilde{\mathbf{d}}$ and threshold peaks higher than $r$ to remove noise.
 $\mathcal{J}$ is then computed between the annotated start-end times and the thresholded peaks.
+8 Repetition Localisation
 VRC metrics only relate predicted to correct counts, regardless of whether the repetitions have been correctly identified.
 We thus investigate whether the peaks of the predicted density map $\tilde{\mathbf{d}}$ align with the annotated start-end times of repetitions in the ground truth.
-Following action localisation methods [8, 18, 22], we adopt and adjust the Jaccard index $\mathcal{J}$ for repetition localisation.
-Using the predicted density map we define different thresholds for salient regions of repetitions.
+Following action localisation methods [8, 20, 25], we adopt the Jaccard index $\mathcal{J}$ for repetition localisation.
 As the values of $\tilde{\mathbf{d}}$ peaks vary across videos, we apply thresholds $\theta$ relative to the maximum and minimum values, $r=\theta(\max(\tilde{\mathbf{d}})-\min(\tilde{\mathbf{d}}))$.
@@
 In contrast, peaks that do not overlap with repetitions are false positives (FP) and repetitions that do not overlap with any peak are false negatives (FN).
-We then calculate the Jaccard index as TP divided by all the correspondences (TP + FP + FN) as customary.
-(a) counts
-(b) repetition length
-(c) video duration
-Figure 10: Comparison of ESCounts and TransRAC [20] for groups of counts/repetition duration/video-length
-In Table 10 we report the Jaccard index over different thresholds alongside the Mean Average Precision (mAP) on RepCount.
+We then calculate $\mathcal{J}$ as TP divided by all the correspondences (TP + FP + FN) as customary.
+In Tab.
+10 we report the Jaccard index over different thresholds alongside the Mean Average Precision (mAP) on RepCount.
 We select TransRAC [22] as a baseline due to their publicly available checkpoint.
-Across thresholds, ESCounts outperforms [20] with the most notable improvements observed over higher threshold values demonstrating that ESCounts predict density maps with higher contrast between higher and lower salient regions.
-For 0.9, 0.8, and 0.7 thresholds ESCounts demonstrates an +8.68%, +10.30%, and +8.10% improvement over [20].
-9 Comparison with TransRAC over Groups
-As in Sec.
-4.3, we divide the videos into groups based on their counts, average repetition duration, and lengths.
-We compare ESCounts’ performance against TransRAC in each of these groups.
-As evident from Fig.
-10, ESCounts performs significantly better than TransRAC on high counts.
-From Fig.
-10, ESCounts performs strongly over varying repetition duration, whilst TransRAC struggles on videos with XS and XL repetitions.
-Fig.
-10 shows that even though both TransRAC and ESCounts performance drops on longer videos, ESCounts perform significantly better on them.
+Across thresholds, ESCounts outperforms [22] with the most notable improvements observed over higher threshold values.
+This demonstrates ESCounts’ ability to predict density maps with higher contrast between higher and lower salient regions.
+For 0.9, 0.8, and 0.7 thresholds ESCounts demonstrates a +8.68%, +10.30%, and +8.10% improvement over [22].
+Table 11: Comparison between ESCounts and TAS baseline on close and open-set RepCount setting.
+Task Method benchmark open-set
+[…]
+TAS GTRM [24] 0.527 0.159 1.000 0.000
+TriDet [53] 0.603 0.232 1.000 0.000
+VRC ESCounts 0.213 0.563 0.436 0.519
+9 Distinction between VRC and TAS
+Unlike Temporal Action Segmentation (TAS) methods, VRC methods can generalise to unseen action classes.
+In Tab.
+11 we compare ESCounts to a TAS method [24] on the RepCount benchmark (close-set) and open-set setting.
+As shown, [24] can only localise the actions of a pre-defined set of categories with which the model was trained.
+In contrast, VRC is learned as an open-set task.
+As ESCounts uses a learnt latent to encode class-independent repetition embeddings, it effectively generalises to unseen categories.
+In addition, ESCounts can better handle large variations in repetition durations that are present in VRC videos compared to [24], which as noted by [22] is a weakness of TAS methods.
+10 Results on egocentric VRC.
+Table 12: Results on OVR-Ego4D.$\dagger$ indicates results have been copied from [16]. (V) corresponds to vision-only models and (V+L) to vision and language models.
+[…]
+V RepNet [15] $\dagger$ 3.20 0.74 0.19 0.43
+ESCounts 2.41 0.32 0.30 0.68
+V+L OVRCounter [16] $\dagger$ 1.60 0.35 0.29 0.66
+The recently-introduced OVR-Ego4D [16] is an Ego4D [19] subset containing clips of repetitive egocentric actions, e.g.
+cutting onions, rolling dough.
+It comprises 50.6K 10-second clips with 41.9K train and 8.7K test clips.
+Annotations are only provided for the number of repetitions and not the individual start and end times per repetition.
+Thus, similar to Countix, we define pseudo-labels to estimate the density maps.
+We evaluate ESCounts on OVR-Ego4D in Tab.
+12.
+Compared to the vision-language-based OVRCounter, [16] ESCounts improves OBZ, OBO, and MAE, with only visual inputs, without any language input in training or inference, showing ESCounts’ effectiveness for the domain of egocentric counting.
+We also add some qualitative results in Fig.
+10.
+Similar to results on other datasets, ESCounts predicts accurate counts a over diverse range of counts.
+The peaks of individual repetitions are not as clear, due to the pseudo-labels, but ESCounts correctly finds the OBO counts in each case.
+Figure 10: Qualitative results of ESCounts on OVR-Ego4D.
+For the selected videos, we show both ground truth (GT) and predicted (P) density maps along with the counts.
+Note that for OVR-Ego4D, we do not have temporal annotations for individual repetitions.
+Therefore similar to Countix, we show pseudo-labels as the GT density maps.